### v2

# Initial value space of the fourth order Painlevé system with $A_{5}^{(1)}\times A_{1}^{(1)}$ symmetry

Abstract
The initial value spaces of the Painlevé equations are proposed by Okamoto. They are symplectic manifolds in which the Painlevé equations are described as polynomial Hamiltonian systems on all coordinates. In this article, we construct an initial value space of the fourth order Painlevé system with affine Weyl group symmetry of type $A_{5}^{(1)}\times A_{1}^{(1)}$.
Key Words: Higher order Painlevé system, Affine Weyl group, Initial value space.
2010 Mathematics Subject Classification: 34M55, 33E17, 37K35.
1 Introduction
The Painlevé equations are nonlinear ordinary differential equations of second order. Each of them is expressed as a polynomial Hamiltonian system. Since it has movable poles, we have to consider compactification of its phase space for the analytic continuation. Then there are infinitely many solutions passing through a point called the accessible singularity in the obtained compact phase space. Hence it becomes the next problem to give a phase space separating those solutions. This problem was solved by Okamoto in [10] with the aid of the blowing up method. The obtained space is now called the initial value space (IVS).
Afterward, Takano and his collaborators gave canonical coordinates for the IVSs of the Painlevé equations explicitly on which the Hamiltonians are polynomials in [5, 7, 15]. They also showed that each Painlevé equation is a unique Hamiltonian system which is holomorphic on all coordinates of the IVS. This property is now called the holomorphy property. Besides, a relationship between the symmetries for the Painlevé equations and the IVSs is pointed out in [8].
Sakai classified the discrete Painlevé equations from viewpoints of symmetries and IVSs in [11]. According to that, both symmetries and IVSs are labeled in terms of the affine root systems and the continuous Painlevé equations can be regarded as continuous flows commuting with some discrete Painlevé equations. On the other hand, several higher order generalizations of the Painlevé equations were proposed in [2, 3, 9, 14, 18, 22]. Hence we expect that there exists a classification theory of them from viewpoints of symmetries and IVSs.
Some IVSs of the higher order Painlevé systems have been investigated as follows.
Painlevé system differential difference (symmetry) holomorphy
Garnier [4], [16] [21] [12]
Noumi-Yamada [19] [1] [14]
Sasano [14]
FST [20] [13]
Kimura constructed the IVS of the Garnier system in $n$ variables in [4] and Suzuki constructed those of the degenerate Garnier systems in $2$ variables in [16]. In both works, the blowing up method is utilized. Tahara constructed the IVS of a fourth order Noumi-Yamada system by using Laurent series expansions of solutions in [19]. Takenawa and Carstea constructed the IVSs of some systems by resolving singularities of symmetries and gave geometric interpretations in terms of the Néron-Severi bilattices. Sasano derived those systems from the holomorphy properties.
In this article, we construct an IVS of the fourth order FST system. This system is derived from a similarity reduction of the Drinfeld-Sokolov hierarchy ([2]) or the UC hierarchy ([22]) and is expressed as the polynomial Hamiltonian system
$\begin{split}&\frac{dq_{i}}{dt}=\frac{\partial H}{\partial p_{i}},\quad\frac{dp_{i}}{dt}=-\frac{\partial H}{\partial q_{i}}\quad(i=1,2),\\ &H=H_{\rm VI}\left[\alpha_{3}+\alpha_{5}-\eta,\alpha_{0},\alpha_{2}+\alpha_{4},\alpha_{1}\eta;q_{1},p_{1}\right]+H_{\rm VI}\left[\alpha_{1}+\alpha_{5}-\eta,\alpha_{0}+\alpha_{2},\alpha_{4},\alpha_{3}\eta;q_{2},p_{2}\right]\\ &\qquad+\frac{(q_{1}-1)(q_{2}-t)\left\{(q_{1}p_{1}+\alpha_{1})p_{2}+(q_{2}p_{2}+\alpha_{3})p_{1}\right\}}{t(t-1)},\end{split}$ (1.1)
where
$H_{\rm{VI}}[a,b,c,d;q,p]=\frac{q(q-1)(q-t)p^{2}-\{a(q-1)(q-t)+bq(q-t)+(c-1)q(q-1)\}p+dq}{t(t-1)}$
and
$\alpha_{0}+\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4}+\alpha_{5}=1.$
Here the independent variable $t$ is defined in $B=\mathbb{C}\setminus\{0,1\}$. The symmetry of system (1.1) was first given in [22]. Afterward, it was shown that the group of the symmetry is isomorphic to an extended affine Weyl group of type $A_{5}^{(1)}\times A_{1}^{(1)}$ in [6]. According to that, this group is generated by the birational transformations $r_{0},\ldots,r_{5},r^{\prime}_{0},\pi_{1},\pi_{2},\rho$. Their actions on the parameters are given as follows.
[…]
Their actions on the dependent and independent variables are given by
$r_{0}(q_{i})=q_{i}\quad(i=1,2),\quad r_{0}(p_{1})=p_{1}-\frac{\alpha_{0}}{q_{1}-1},\quad r_{0}(p_{2})=p_{2},\quad r_{0}(t)=t,$
$r_{1}(q_{1})=q_{1}+\frac{\alpha_{1}}{p_{1}},\quad r_{1}(q_{2})=q_{2},\quad r_{1}(p_{i})=p_{i}\quad(i=1,2),\quad r_{1}(t)=t,$
$r_{2}(q_{i})=q_{i}\quad(i=1,2),\quad r_{2}(p_{1})=p_{1}-\frac{\alpha_{2}}{q_{1}-q_{2}},\quad r_{2}(p_{2})=p_{2}+\frac{\alpha_{2}}{q_{1}-q_{2}},\quad r_{2}(t)=t,$
$r_{3}(q_{1})=q_{1},\quad r_{3}(q_{2})=q_{2}+\frac{\alpha_{3}}{p_{2}},\quad r_{3}(p_{i})=p_{i}\quad(i=1,2),\quad r_{3}(t)=t,$
$r_{4}(q_{i})=q_{i}\quad(i=1,2),\quad r_{4}(p_{1})=p_{1},\quad r_{4}(p_{2})=p_{2}-\frac{\alpha_{4}}{q_{2}-t},\quad r_{4}(t)=t,$
$r_{5}(q_{i})=q_{i}\frac{q_{1}p_{1}+q_{2}p_{2}+\eta}{q_{1}p_{1}+q_{2}p_{2}-\alpha_{5}+\eta},\quad r_{5}(p_{i})=p_{i}\frac{q_{1}p_{1}+q_{2}p_{2}-\alpha_{5}+\eta}{q_{1}p_{1}+q_{2}p_{2}+\eta}\quad(i=1,2),\quad r_{5}(t)=t,$
$r^{\prime}_{0}(q_{i})=\frac{1}{q_{i}},\quad r^{\prime}_{0}(p_{i})=-q_{i}(q_{i}p_{i}+\alpha_{2i-1})\quad(i=1,2),\quad r^{\prime}_{0}(t)=\frac{1}{t},$
$\pi_{1}(q_{1})=\frac{(q_{1}-t)p_{1}+(q_{2}-t)p_{2}+\eta}{(q_{1}-1)p_{1}+(q_{2}-t)p_{2}+\eta},\quad\pi_{1}(q_{2})=\frac{(q_{1}-t)p_{1}+(q_{2}-t)p_{2}+\eta}{(q_{1}-1)p_{1}+(q_{2}-1)p_{2}+\eta},$
$\pi_{1}(q_{1}p_{1})=\frac{q_{1}-q_{2}}{t-1}\left\{(q_{1}-1)p_{1}+(q_{2}-t)p_{2}+\eta\right\}-\alpha_{2},$
$\pi_{1}(q_{2}p_{2})=\frac{q_{2}-t}{t-1}\{(q_{1}-1)p_{1}+(q_{2}-1)p_{2}+\eta\}-\alpha_{4},\quad\pi_{1}(t)=t,$
$\pi_{2}(q_{1})=\frac{q_{2}}{q_{1}},\quad\pi_{2}(q_{2})=\frac{t}{q_{1}},\quad\pi_{2}(p_{1})=q_{1}p_{2},\quad\pi_{2}(p_{2})=-\frac{q_{1}(q_{1}p_{1}+q_{2}p_{2}+\eta)}{t},\quad\pi_{2}(t)=t,$
$\rho(q_{1})=q_{1},\quad\rho(q_{2})=\frac{tq_{1}}{q_{2}},\quad\rho(p_{1})=\frac{q_{1}p_{1}+q_{2}p_{2}-\alpha_{5}+\eta}{q_{1}},\quad\rho(p_{2})=-\frac{q_{2}(q_{2}p_{2}+\alpha_{3})}{tq_{1}},\quad\rho(t)=t.$
System (1.1) is invariant under their actions. Moreover, groups $\langle r_{0},\ldots,r_{5}\rangle$ and $\langle r^{\prime}_{0},r^{\prime}_{0}\pi_{2}^{-1}\pi_{1}^{2}\rangle$ are isomorphic to the affine Weyl group of type $A_{5}^{(1)}$ and $A_{1}^{(1)}$ respectively.
This article is organized as follows. In Section 2, we construct a compactification of the phase space of system (1.1) called the Hirzebruch manifold following the previous work [4, 16]. In Section 3, we determine the accessible singularities in the Hirzebruch manifold. As a result, two types of the accessible singularities $C_{j}^{(kl)}$ and $P_{j}^{(kl)}$ are obtained. In Section 4, we resolve $C_{j}^{(kl)}$ in two ways, the blowing up method and utilizing formal power series. In Section 5, we resolve $P_{j}^{(kl)}$ by using formal power series and the symmetry. In Section 6, we state the main result and consider the holomorphy property.
2 Compactification of the initial phase space
In this section, we construct a four dimensional Hirzebruch manifold $\Sigma_{\eta}$ as a compactification of the phase space $\mathbb{C}^{4}$ of system (1.1) following the previous work [4, 16].
Let $\xi$ and $\zeta^{(i)}\ (i=0,1,2)$ be homogeneous coordinates of $\mathbb{P}^{2}$ defined by
$\xi=\left[\xi_{0}:\xi_{1}:\xi_{2}\right],\quad\zeta^{(i)}=\left[\zeta_{0}^{(i)}:\zeta_{1}^{(i)}:\zeta_{2}^{(i)}\right]$
with relations
[…]
We also set
$W_{ij}=\left\{\left(\xi,\zeta^{(i)}\right)\in\mathbb{P}^{2}\times\mathbb{P}^{2}\bigm{|}\xi_{i}\neq 0,\ \zeta_{j}^{(i)}\neq 0\right\},\quad W_{i}=W_{i0}\cup W_{i1}\cup W_{i2}\quad(i,j=0,1,2).$
Since $\det Z_{10}=-\xi_{0}^{3}\xi_{1}^{3}$ and $\det Z_{20}=-\xi_{0}^{3}\xi_{2}^{3}$, we can define manifolds $\Sigma_{\eta}^{0}$ and $\Sigma_{\eta}$ to be the quotient spaces of $W_{00}\cup W_{10}\cup W_{20}$ and $W_{0}\cup W_{1}\cup W_{2}$ respectively by relation (2.1). Then we take affine coordinates of $W_{ij}$ as follows.
$\begin{split}&W_{00}:(q_{1},q_{2},p_{1},p_{2})=\left(\frac{\xi_{1}}{\xi_{0}},\frac{\xi_{2}}{\xi_{0}},\frac{\zeta_{1}^{(0)}}{\zeta_{0}^{(0)}},\frac{\zeta_{2}^{(0)}}{\zeta_{0}^{(0)}}\right).\\ &W_{01}:\left(q_{1},q_{2},p_{1}^{(01)},p_{2}^{(01)}\right)=\left(\frac{\xi_{1}}{\xi_{0}},\frac{\xi_{2}}{\xi_{0}},\frac{\zeta_{0}^{(0)}}{\zeta_{1}^{(0)}},\frac{\zeta_{2}^{(0)}}{\zeta_{1}^{(0)}}\right).\\ &W_{02}:\left(q_{1},q_{2},p_{1}^{(02)},p_{2}^{(02)}\right)=\left(\frac{\xi_{1}}{\xi_{0}},\frac{\xi_{2}}{\xi_{0}},\frac{\zeta_{1}^{(0)}}{\zeta_{2}^{(0)}},\frac{\zeta_{0}^{(0)}}{\zeta_{2}^{(0)}}\right).\\ &W_{10}:\left(q_{1}^{(1)},q_{2}^{(1)},p_{1}^{(1)},p_{2}^{(1)}\right)=\left(\frac{\xi_{0}}{\xi_{1}},\frac{\xi_{2}}{\xi_{1}},\frac{\zeta_{1}^{(1)}}{\zeta_{0}^{(1)}},\frac{\zeta_{2}^{(1)}}{\zeta_{0}^{(1)}}\right).\\ &W_{11}:\left(q_{1}^{(1)},q_{2}^{(1)},p_{1}^{(11)},p_{2}^{(11)}\right)=\left(\frac{\xi_{0}}{\xi_{1}},\frac{\xi_{2}}{\xi_{1}},\frac{\zeta_{0}^{(1)}}{\zeta_{1}^{(1)}},\frac{\zeta_{2}^{(1)}}{\zeta_{1}^{(1)}}\right).\\ &W_{12}:\left(q_{1}^{(1)},q_{2}^{(1)},p_{1}^{(12)},p_{2}^{(12)}\right)=\left(\frac{\xi_{0}}{\xi_{1}},\frac{\xi_{2}}{\xi_{1}},\frac{\zeta_{1}^{(1)}}{\zeta_{2}^{(1)}},\frac{\zeta_{0}^{(1)}}{\zeta_{2}^{(1)}}\right).\\ &W_{20}:\left(q_{1}^{(2)},q_{2}^{(2)},p_{1}^{(2)},p_{2}^{(2)}\right)=\left(\frac{\xi_{1}}{\xi_{2}},\frac{\xi_{0}}{\xi_{2}},\frac{\zeta_{1}^{(2)}}{\zeta_{0}^{(2)}},\frac{\zeta_{2}^{(2)}}{\zeta_{0}^{(2)}}\right).\\ &W_{21}:\left(q_{1}^{(2)},q_{2}^{(2)},p_{1}^{(21)},p_{2}^{(21)}\right)=\left(\frac{\xi_{1}}{\xi_{2}},\frac{\xi_{0}}{\xi_{2}},\frac{\zeta_{0}^{(2)}}{\zeta_{1}^{(2)}},\frac{\zeta_{2}^{(2)}}{\zeta_{1}^{(2)}}\right).\\ &W_{22}:\left(q_{1}^{(2)},q_{2}^{(2)},p_{1}^{(22)},p_{2}^{(22)}\right)=\left(\frac{\xi_{1}}{\xi_{2}},\frac{\xi_{0}}{\xi_{2}},\frac{\zeta_{1}^{(2)}}{\zeta_{2}^{(2)}},\frac{\zeta_{0}^{(2)}}{\zeta_{2}^{(2)}}\right).\end{split}$ (2.2)
It implies relations
$\begin{split}&\left(q_{1}^{(1)},q_{2}^{(1)},p_{1}^{(1)},p_{2}^{(1)}\right)=\left(\frac{1}{q_{1}},\frac{q_{2}}{q_{1}},-q_{1}(q_{1}p_{1}+q_{2}p_{2}+\eta),q_{1}p_{2}\right),\\ &\left(q_{1}^{(2)},q_{2}^{(2)},p_{1}^{(2)},p_{2}^{(2)}\right)=\left(\frac{q_{1}}{q_{2}},\frac{1}{q_{2}},q_{2}p_{1},-q_{2}(q_{1}p_{1}+q_{2}p_{2}+\eta)\right)\end{split}$ (2.3)
and
$\begin{split}&\left(p_{1}^{(01)},p_{2}^{(01)}\right)=\left(\frac{1}{p_{1}},\frac{p_{2}}{p_{1}}\right),\quad\left(p_{1}^{(02)},p_{2}^{(02)}\right)=\left(\frac{p_{1}}{p_{2}},\frac{1}{p_{2}}\right),\\ &\left(p_{1}^{(11)},p_{2}^{(11)}\right)=\left(\frac{1}{p_{1}^{(1)}},\frac{p_{2}^{(1)}}{p_{1}^{(1)}}\right),\quad\left(p_{1}^{(12)},p_{2}^{(12)}\right)=\left(\frac{p_{1}^{(1)}}{p_{2}^{(1)}},\frac{1}{p_{2}^{(1)}}\right),\\ &\left(p_{1}^{(21)},p_{2}^{(21)}\right)=\left(\frac{1}{p_{1}^{(2)}},\frac{p_{2}^{(2)}}{p_{1}^{(2)}}\right),\quad\left(p_{1}^{(22)},p_{2}^{(22)}\right)=\left(\frac{p_{1}^{(2)}}{p_{2}^{(2)}},\frac{1}{p_{2}^{(2)}}\right).\end{split}$ (2.4)
Note that $W_{ij}\simeq\mathbb{C}^{4}$ and $W_{i}\simeq\mathbb{C}^{2}\times\mathbb{P}^{2}$. Also note that $\Sigma_{\eta}\simeq T^{*}\mathbb{P}^{2}\sqcup\left(\mathbb{P}^{2}\times\mathbb{P}^{1}\right)$ if $\eta=0$ and $\Sigma_{\eta}\simeq\mathbb{P}^{2}\times\mathbb{P}^{2}$ if $\eta\neq 0$; see [4, 16] for its detail.
We extend system (1.1) to that on $\Sigma_{\eta}\times B$ (or resp. $\Sigma_{\eta}^{0}\times B$) and denote it by $\mathcal{H}$ (or resp. $\mathcal{H}^{0}$).
Proposition 2.1.
The system $\mathcal{H}^{0}$ is described as the polynomial Hamiltonian system on any affine chart $W_{i0}$.
Proof.
Let $\Psi_{1},\Psi_{2}$ be transformations which act on the initial coordinates $q_{1},q_{2},p_{1},p_{2}$ as
$\Psi_{1}((q_{1},q_{2},p_{1},p_{2}))=\left(q_{2},tq_{1},p_{2},\frac{p_{1}}{t}\right),\quad\Psi_{2}((q_{1},q_{2},p_{1},p_{2}))=\left(tq_{2},tq_{1},\frac{p_{2}}{t},\frac{p_{1}}{t}\right).$
Then we obtain
$\left(q_{1}^{(1)},q_{2}^{(1)},p_{1}^{(1)},p_{2}^{(1)}\right)=\Psi_{1}\pi_{2}^{-1}((q_{1},q_{2},p_{1},p_{2})),\quad\left(q_{1}^{(2)},q_{2}^{(2)},p_{1}^{(2)},p_{2}^{(2)}\right)=\Psi_{2}\pi_{2}((q_{1},q_{2},p_{1},p_{2})).$
Since the transformation $\pi_{2}$ provides the symmetry of system (1.1), the system on $W_{10}\times B$ and $W_{20}\times B$ are both polynomial Hamiltonian systems. Note that both relations in (2.3) are symplectic transformations. ∎
Proposition 2.2.
If we assume that
$\alpha_{1}-\eta\neq 0,\quad\alpha_{3}-\eta\neq 0,\quad\alpha_{1}+\alpha_{3}-\eta\neq 0,\quad\alpha_{1}+\alpha_{2}+\alpha_{3}-\eta\neq 0,$ (2.5)
any solution to the system $\mathcal{H}^{0}$ passes through $W_{00}\times B$.
Proof.
We assume that there exists a solution to the system $\mathcal{H}^{0}$ satisfying the condition $\xi_{0}\equiv 0$, namely $q_{1}^{(1)}\equiv 0$ or $q_{2}^{(2)}\equiv 0$.
We first consider the condition $q_{1}^{(1)}\equiv 0$. Then the equation for $q_{1}^{(1)}$ in $\mathcal{H}^{0}$ is described as
$\frac{dq_{1}^{(1)}}{dt}=-\frac{\left(q_{2}^{(1)}\right)^{2}p_{2}^{(1)}-q_{2}^{(1)}p_{2}^{(1)}+\alpha_{3}q_{2}^{(1)}+\alpha_{1}-\eta}{t(t-1)}.$
We set
$Q_{1}=\left(q_{2}^{(1)}\right)^{2}p_{2}^{(1)}-q_{2}^{(1)}p_{2}^{(1)}+\alpha_{3}q_{2}^{(1)}+\alpha_{1}-\eta.$
Then we obtain $Q_{1}\equiv 0$, which implies $q_{2}^{(1)}\not\equiv 0,1$ due to assumption (2.5). Moreover, the equation for $Q_{1}$ in $\mathcal{H}^{0}$ is described as
$\frac{dQ_{1}}{dt}=-\frac{(\alpha_{1}+\alpha_{3}-\eta)(\alpha_{1}+\alpha_{2}+\alpha_{3}-\eta)q_{2}^{(1)}}{t\left(q_{2}^{(1)}-1\right)}.$
The equality $(\alpha_{1}+\alpha_{3}-\eta)(\alpha_{1}+\alpha_{2}+\alpha_{3}-\eta)q_{2}^{(1)}\equiv 0$ contradicts with $q_{2}^{(1)}\not\equiv 0$ and (2.5). Therefore there does not exist any solution satisfying $q_{1}^{(1)}\equiv 0$.
We can prove for $q_{2}^{(2)}\equiv 0$ in a similar manner. The second condition of (2.5) is required in this case. ∎
Remark 2.3.
Under the condition $\alpha_{1}-\eta=0$, the system $\mathcal{H}^{0}$ on $W_{10}\times B$ admits a particular solution such that $q_{1}^{(1)}=q_{2}^{(1)}=0$. Then the equations for $p_{1}^{(1)},p_{2}^{(1)}$ are described as
$t(1-t)\frac{dp_{1}^{(1)}}{dt}=\left(p_{1}^{(1)}\right)^{2}+\{-(\alpha_{0}+\alpha_{1}+\alpha_{5})t+\alpha_{0}\}p_{1}^{(1)}+\alpha_{1}\alpha_{5}t+\left(p_{1}^{(1)}-\alpha_{5}\right)tp_{2}^{(1)},$
$t(1-t)\frac{dp_{2}^{(1)}}{dt}=t\left(p_{2}^{(1)}\right)^{2}+\{\alpha_{2}t-(\alpha_{1}+\alpha_{2}+\alpha_{3})\}p_{2}^{(1)}+\alpha_{1}\alpha_{3}+\left(p_{2}^{(1)}-\alpha_{3}\right)p_{1}^{(1)},$
whose solutions are given as ratios of the generalized hypergeometric series $_{3}F_{2}$; see [17] for its detail. In a similar manner, we can derive particular solutions for $\alpha_{3}-\eta=0$, $\alpha_{1}+\alpha_{3}-\eta=0$ and $\alpha_{1}+\alpha_{2}+\alpha_{3}-\eta=0$.
Proposition 2.4.
The system $\mathcal{H}$ has pole singularities on $\left(\Sigma_{\eta}\setminus\Sigma_{\eta}^{0}\right)\times B$.
Proof.
We first investigate the singularities of the system $\mathcal{H}$ on $W_{0}\times B$. The system $\mathcal{H}$ on $W_{01}\times B$ is described as
$\frac{dq_{1}}{dt}=\frac{F_{1}^{(01)}}{p_{1}^{(01)}},\quad\frac{dq_{2}}{dt}=\frac{F_{2}^{(01)}}{p_{1}^{(01)}},\quad\frac{dp_{1}^{(01)}}{dt}=G_{1}^{(01)},\quad\frac{dp_{2}^{(01)}}{dt}=\frac{G_{2}^{(01)}}{p_{1}^{(01)}},$ (2.6)
where $F_{1}^{(01)},F_{2}^{(01)},G_{1}^{(01)},G_{2}^{(01)}\in\mathbb{C}(t)\left[q_{1},q_{2},p_{1}^{(01)},p_{2}^{(01)}\right]$. Since the numerators $F_{1}^{(01)},F_{2}^{(01)},G_{2}^{(01)}$ are not divisible by $p_{1}^{(01)}$, the singularity of system (2.6) is
$\left\{\left(q_{1},q_{2},p_{1}^{(01)},p_{2}^{(01)}\right)\in\mathbb{C}^{4}\bigm{|}p_{1}^{(01)}=0\right\}\times B.$
Similarly, the singularity of the system $\mathcal{H}$ on $W_{02}\times B$ is
$\left\{\left(q_{1},q_{2},p_{1}^{(02)},p_{2}^{(02)}\right)\in\mathbb{C}^{4}\bigm{|}p_{2}^{(02)}=0\right\}\times B.$
By using $p_{1}^{(01)}=\frac{\zeta_{0}^{(0)}}{\zeta_{1}^{(0)}}$ and $p_{2}^{(02)}=\frac{\zeta_{0}^{(0)}}{\zeta_{2}^{(0)}}$, we can rewrite the singularity on $W_{0}\times B$ in terms of the homogeneous coordinate (2.2) as
$\left\{\left(\xi,\zeta^{(0)}\right)\in\mathbb{P}^{2}\times\mathbb{P}^{2}\bigm{|}\xi_{0}\neq 0,\ \zeta_{0}^{(0)}=0\right\}\times B.$
Note that the system $\mathcal{H}$ coincides with system (1.1) and has no singularity on $W_{00}\times B$.
We can obtain the singularities of the system $\mathcal{H}$ on $W_{1}\times B$ and $W_{2}\times B$ as
$\left\{\left(\xi,\zeta^{(1)}\right)\in\mathbb{P}^{2}\times\mathbb{P}^{2}\bigm{|}\xi_{1}\neq 0,\ \zeta_{0}^{(1)}=0\right\}\times B,\quad\left\{\left(\xi,\zeta^{(2)}\right)\in\mathbb{P}^{2}\times\mathbb{P}^{2}\bigm{|}\xi_{2}\neq 0,\ \zeta_{0}^{(2)}=0\right\}\times B$
in a similar manner. Hence the system $\mathcal{H}$ has pole singularities on $\left(\Sigma_{\eta}\setminus\Sigma_{\eta}^{0}\right)\times B$. ∎
3 Accessible singularities
In the previous section, we have shown that the system $\mathcal{H}$ has the singularities only on $W_{ij}\times B\ (i=0,1,2,\ j=1,2)$. On these $W_{ij}\times B$, there exist accessible singularities where solutions to the system $\mathcal{H}^{0}$ pass through. For example, the accessible singularities on $W_{01}\times B$ are given by the condition $p_{1}^{(01)}=F_{1}^{(01)}=F_{2}^{(01)}=G_{2}^{(01)}=0$. It follows from the form of system (2.6). In this section, we list all of the accessible singularities of the system $\mathcal{H}$.
We first give the following lemma.
Lemma 3.1.
All of the accessible singularities of the system $\mathcal{H}$ are given by
$C_{1}^{(01)}=\left\{\left(q_{1},q_{2},p_{1}^{(01)},p_{2}^{(01)},t\right)\bigm{|}q_{1}-q_{2}=p_{1}^{(01)}=p_{2}^{(01)}+1=0\right\},$
$C_{2}^{(01)}=\left\{\left(q_{1},q_{2},p_{1}^{(01)},p_{2}^{(01)},t\right)\bigm{|}q_{1}-1=p_{1}^{(01)}=p_{2}^{(01)}=0\right\},$
$C_{3}^{(01)}=\left\{\left(q_{1},q_{2},p_{1}^{(01)},p_{2}^{(01)},t\right)\bigm{|}q_{1}=q_{2}=p_{1}^{(01)}=0\right\},$
$P_{1}^{(01)}=\left\{\left(q_{1},q_{2},p_{1}^{(01)},p_{2}^{(01)},t\right)\bigm{|}q_{1}-t=q_{2}-t=p_{1}^{(01)}=p_{2}^{(01)}=0\right\},$
$P_{2}^{(01)}=\left\{\left(q_{1},q_{2},p_{1}^{(01)},p_{2}^{(01)},t\right)\bigm{|}q_{1}-t=q_{2}+t=p_{1}^{(01)}=p_{2}^{(01)}=0\right\},$
$P_{3}^{(01)}=\left\{\left(q_{1},q_{2},p_{1}^{(01)},p_{2}^{(01)},t\right)\bigm{|}q_{1}=q_{2}-t=p_{1}^{(01)}=p_{2}^{(01)}=0\right\},$
$P_{4}^{(01)}=\left\{\left(q_{1},q_{2},p_{1}^{(01)},p_{2}^{(01)},t\right)\bigm{|}q_{1}=q_{2}-t=p_{1}^{(01)}=(t-1)p_{2}^{(01)}-2=0\right\},$
$P_{5}^{(01)}=\left\{\left(q_{1},q_{2},p_{1}^{(01)},p_{2}^{(01)},t\right)\bigm{|}q_{1}-1=q_{2}=p_{1}^{(01)}=2tp_{2}^{(01)}+t-1=0\right\},$
$P_{6}^{(01)}=\left\{\left(q_{1},q_{2},p_{1}^{(01)},p_{2}^{(01)},t\right)\bigm{|}q_{1}-1=q_{2}-t=p_{1}^{(01)}=tp_{2}^{(01)}+1=0\right\},$
$P_{7}^{(01)}=\left\{\left(q_{1},q_{2},p_{1}^{(01)},p_{2}^{(01)},t\right)\bigm{|}q_{1}+1=q_{2}+t=p_{1}^{(01)}=tp_{2}^{(01)}+1=0\right\}$
on $W_{01}\times B$,
$C_{1}^{(02)}=\left\{\left(q_{1},q_{2},p_{1}^{(02)},p_{2}^{(02)},t\right)\bigm{|}q_{1}-q_{2}=p_{1}^{(02)}+1=p_{2}^{(02)}=0\right\},$
$C_{2}^{(02)}=\left\{\left(q_{1},q_{2},p_{1}^{(02)},p_{2}^{(02)},t\right)\bigm{|}q_{2}-t=p_{1}^{(02)}=p_{2}^{(02)}=0\right\},$
$C_{3}^{(02)}=\left\{\left(q_{1},q_{2},p_{1}^{(02)},p_{2}^{(02)},t\right)\bigm{|}q_{1}=q_{2}=p_{2}^{(02)}=0\right\},$
$P_{1}^{(02)}=\left\{\left(q_{1},q_{2},p_{1}^{(02)},p_{2}^{(02)},t\right)\bigm{|}q_{1}-1=q_{2}-1=p_{1}^{(02)}=p_{2}^{(02)}=0\right\},$
$P_{2}^{(02)}=\left\{\left(q_{1},q_{2},p_{1}^{(02)},p_{2}^{(02)},t\right)\bigm{|}q_{1}+1=q_{2}-1=p_{1}^{(02)}=p_{2}^{(02)}=0\right\},$
$P_{3}^{(02)}=\left\{\left(q_{1},q_{2},p_{1}^{(02)},p_{2}^{(02)},t\right)\bigm{|}q_{1}-1=q_{2}=p_{1}^{(02)}=p_{2}^{(02)}=0\right\},$
$P_{4}^{(02)}=\left\{\left(q_{1},q_{2},p_{1}^{(02)},p_{2}^{(02)},t\right)\bigm{|}q_{1}=q_{2}-t=2p_{1}^{(02)}-t+1=p_{2}^{(02)}=0\right\},$
$P_{5}^{(02)}=\left\{\left(q_{1},q_{2},p_{1}^{(02)},p_{2}^{(02)},t\right)\bigm{|}q_{1}-1=q_{2}=(t-1)p_{1}^{(02)}+2t=p_{2}^{(02)}=0\right\},$
$P_{6}^{(02)}=\left\{\left(q_{1},q_{2},p_{1}^{(02)},p_{2}^{(02)},t\right)\bigm{|}q_{1}-1=q_{2}-t=p_{1}^{(02)}+t=p_{2}^{(02)}=0\right\},$
$P_{7}^{(02)}=\left\{\left(q_{1},q_{2},p_{1}^{(02)},p_{2}^{(02)},t\right)\bigm{|}q_{1}+1=q_{2}+t=p_{1}^{(02)}+t=p_{2}^{(02)}=0\right\}$
on $W_{02}\times B$,
$C_{1}^{(11)}=\left\{\left(q_{1}^{(1)},q_{2}^{(1)},p_{1}^{(11)},p_{2}^{(11)},t\right)\bigm{|}tq_{1}^{(1)}-q_{2}^{(1)}=p_{1}^{(11)}=tp_{2}^{(11)}+1=0\right\},$
$C_{2}^{(11)}=\left\{\left(q_{1}^{(1)},q_{2}^{(1)},p_{1}^{(11)},p_{2}^{(11)},t\right)\bigm{|}q_{1}^{(1)}-1=p_{1}^{(11)}=p_{2}^{(11)}=0\right\},$
$C_{3}^{(11)}=\left\{\left(q_{1}^{(1)},q_{2}^{(1)},p_{1}^{(11)},p_{2}^{(11)},t\right)\bigm{|}q_{1}^{(1)}=q_{2}^{(1)}=p_{1}^{(11)}=0\right\},$
$P_{1}^{(11)}=\left\{\left(q_{1}^{(1)},q_{2}^{(1)},p_{1}^{(11)},p_{2}^{(11)},t\right)\bigm{|}tq_{1}^{(1)}-1=q_{2}^{(1)}-1=p_{1}^{(11)}=p_{2}^{(11)}=0\right\},$
$P_{2}^{(11)}=\left\{\left(q_{1}^{(1)},q_{2}^{(1)},p_{1}^{(11)},p_{2}^{(11)},t\right)\bigm{|}tq_{1}^{(1)}-1=q_{2}^{(1)}+1=p_{1}^{(11)}=p_{2}^{(11)}=0\right\},$
$P_{3}^{(11)}=\left\{\left(q_{1}^{(1)},q_{2}^{(1)},p_{1}^{(11)},p_{2}^{(11)},t\right)\bigm{|}q_{1}^{(1)}=q_{2}^{(1)}-1=p_{1}^{(11)}=p_{2}^{(11)}=0\right\},$
$P_{4}^{(11)}=\left\{\left(q_{1}^{(1)},q_{2}^{(1)},p_{1}^{(11)},p_{2}^{(11)},t\right)\bigm{|}q_{1}^{(1)}-1=q_{2}^{(1)}=p_{1}^{(11)}=2tp_{2}^{(11)}-t+1=0\right\},$
$P_{5}^{(11)}=\left\{\left(q_{1}^{(1)},q_{2}^{(1)},p_{1}^{(11)},p_{2}^{(11)},t\right)\bigm{|}q_{1}^{(1)}=q_{2}^{(1)}-1=p_{1}^{(11)}=(t-1)p_{2}^{(11)}+2=0\right\},$
$P_{6}^{(11)}=\left\{\left(q_{1}^{(1)},q_{2}^{(1)},p_{1}^{(11)},p_{2}^{(11)},t\right)\bigm{|}q_{1}^{(1)}-1=q_{2}^{(1)}-1=p_{1}^{(11)}=p_{2}^{(11)}+1=0\right\},$
$P_{7}^{(11)}=\left\{\left(q_{1}^{(1)},q_{2}^{(1)},p_{1}^{(11)},p_{2}^{(11)},t\right)\bigm{|}q_{1}^{(1)}+1=q_{2}^{(1)}+1=p_{1}^{(11)}=p_{2}^{(11)}+1=0\right\}$
on $W_{11}\times B$,
$C_{1}^{(12)}=\left\{\left(q_{1}^{(1)},q_{2}^{(1)},p_{1}^{(12)},p_{2}^{(12)},t\right)\bigm{|}tq_{1}^{(1)}-q_{2}^{(1)}=p_{1}^{(12)}+t=p_{2}^{(12)}=0\right\},$
$C_{2}^{(12)}=\left\{\left(q_{1}^{(1)},q_{2}^{(1)},p_{1}^{(12)},p_{2}^{(12)},t\right)\bigm{|}q_{2}^{(1)}-1=p_{1}^{(12)}=p_{2}^{(12)}=0\right\},$
$C_{3}^{(12)}=\left\{\left(q_{1}^{(1)},q_{2}^{(1)},p_{1}^{(12)},p_{2}^{(12)},t\right)\bigm{|}q_{1}^{(1)}=q_{2}^{(1)}=p_{2}^{(12)}=0\right\},$
$P_{1}^{(12)}=\left\{\left(q_{1}^{(1)},q_{2}^{(1)},p_{1}^{(12)},p_{2}^{(12)},t\right)\bigm{|}q_{1}^{(1)}-1=q_{2}^{(1)}-t=p_{1}^{(12)}=p_{2}^{(12)}=0\right\},$
$P_{2}^{(12)}=\left\{\left(q_{1}^{(1)},q_{2}^{(1)},p_{1}^{(12)},p_{2}^{(12)},t\right)\bigm{|}q_{1}^{(1)}+1=q_{2}^{(1)}-t=p_{1}^{(12)}=p_{2}^{(12)}=0\right\},$
$P_{3}^{(12)}=\left\{\left(q_{1}^{(1)},q_{2}^{(1)},p_{1}^{(12)},p_{2}^{(12)},t\right)\bigm{|}q_{1}^{(1)}-1=q_{2}^{(1)}=p_{1}^{(12)}=p_{2}^{(12)}=0\right\},$
$P_{4}^{(12)}=\left\{\left(q_{1}^{(1)},q_{2}^{(1)},p_{1}^{(12)},p_{2}^{(12)},t\right)\bigm{|}q_{1}^{(1)}-1=q_{2}^{(1)}=(t-1)p_{1}^{(12)}-2t=p_{2}^{(12)}=0\right\},$
$P_{5}^{(12)}=\left\{\left(q_{1}^{(1)},q_{2}^{(1)},p_{1}^{(12)},p_{2}^{(12)},t\right)\bigm{|}q_{1}^{(1)}=q_{2}^{(1)}-1=2p_{1}^{(12)}+t-1=p_{2}^{(12)}=0\right\},$
$P_{6}^{(12)}=\left\{\left(q_{1}^{(1)},q_{2}^{(1)},p_{1}^{(12)},p_{2}^{(12)},t\right)\bigm{|}q_{1}^{(1)}-1=q_{2}^{(1)}-1=p_{1}^{(12)}+1=p_{2}^{(12)}=0\right\},$
$P_{7}^{(12)}=\left\{\left(q_{1}^{(1)},q_{2}^{(1)},p_{1}^{(12)},p_{2}^{(12)},t\right)\bigm{|}q_{1}^{(1)}+1=q_{2}^{(1)}+1=p_{1}^{(12)}+1=p_{2}^{(12)}=0\right\}$
on $W_{12}\times B$,
$C_{1}^{(21)}=\left\{\left(q_{1}^{(2)},q_{2}^{(2)},p_{1}^{(21)},p_{2}^{(21)},t\right)\bigm{|}q_{1}^{(2)}-q_{2}^{(2)}=p_{1}^{(21)}=p_{2}^{(21)}+1=0\right\},$
$C_{2}^{(21)}=\left\{\left(q_{1}^{(2)},q_{2}^{(2)},p_{1}^{(21)},p_{2}^{(21)},t\right)\bigm{|}q_{1}^{(2)}-1=p_{1}^{(21)}=p_{2}^{(21)}=0\right\},$
$C_{3}^{(21)}=\left\{\left(q_{1}^{(2)},q_{2}^{(2)},p_{1}^{(21)},p_{2}^{(21)},t\right)\bigm{|}q_{1}^{(2)}=q_{2}^{(2)}=p_{1}^{(21)}=0\right\},$
$P_{1}^{(21)}=\left\{\left(q_{1}^{(2)},q_{2}^{(2)},p_{1}^{(21)},p_{2}^{(21)},t\right)\bigm{|}tq_{1}^{(2)}-1=tq_{2}^{(2)}-1=p_{1}^{(21)}=p_{2}^{(21)}=0\right\},$
$P_{2}^{(21)}=\left\{\left(q_{1}^{(2)},q_{2}^{(2)},p_{1}^{(21)},p_{2}^{(21)},t\right)\bigm{|}tq_{1}^{(2)}-1=tq_{2}^{(2)}+1=p_{1}^{(21)}=p_{2}^{(21)}=0\right\},$
$P_{3}^{(21)}=\left\{\left(q_{1}^{(2)},q_{2}^{(2)},p_{1}^{(21)},p_{2}^{(21)},t\right)\bigm{|}q_{1}^{(2)}=tq_{2}^{(2)}-1=p_{1}^{(21)}=p_{2}^{(21)}=0\right\},$
$P_{4}^{(21)}=\left\{\left(q_{1}^{(2)},q_{2}^{(2)},p_{1}^{(21)},p_{2}^{(21)},t\right)\bigm{|}q_{1}^{(2)}-1=q_{2}^{(2)}=p_{1}^{(21)}=2p_{2}^{(21)}-t+1=0\right\},$
$P_{5}^{(21)}=\left\{\left(q_{1}^{(2)},q_{2}^{(2)},p_{1}^{(21)},p_{2}^{(21)},t\right)\bigm{|}q_{1}^{(2)}=tq_{2}^{(2)}-1=p_{1}^{(21)}=(t-1)p_{2}^{(21)}+2t=0\right\},$
$P_{6}^{(21)}=\left\{\left(q_{1}^{(2)},q_{2}^{(2)},p_{1}^{(21)},p_{2}^{(21)},t\right)\bigm{|}q_{1}^{(2)}-1=tq_{2}^{(2)}-1=p_{1}^{(21)}=p_{2}^{(21)}+t=0\right\},$
$P_{7}^{(21)}=\left\{\left(q_{1}^{(2)},q_{2}^{(2)},p_{1}^{(21)},p_{2}^{(21)},t\right)\bigm{|}q_{1}^{(2)}+1=tq_{2}^{(2)}+1=p_{1}^{(21)}=p_{2}^{(21)}+t=0\right\}$
on $W_{21}\times B$ and
$C_{1}^{(22)}=\left\{\left(q_{1}^{(2)},q_{2}^{(2)},p_{1}^{(22)},p_{2}^{(22)},t\right)\bigm{|}q_{1}^{(2)}-q_{2}^{(2)}=p_{1}^{(22)}+1=p_{2}^{(22)}=0\right\},$
$C_{2}^{(22)}=\left\{\left(q_{1}^{(2)},q_{2}^{(2)},p_{1}^{(22)},p_{2}^{(22)},t\right)\bigm{|}tq_{2}^{(2)}-1=p_{1}^{(22)}=p_{2}^{(22)}=0\right\},$
$C_{3}^{(22)}=\left\{\left(q_{1}^{(2)},q_{2}^{(2)},p_{1}^{(22)},p_{2}^{(22)},t\right)\bigm{|}q_{1}^{(2)}=q_{2}^{(2)}=p_{2}^{(22)}=0\right\},$
$P_{1}^{(22)}=\left\{\left(q_{1}^{(2)},q_{2}^{(2)},p_{1}^{(22)},p_{2}^{(22)},t\right)\bigm{|}q_{1}^{(2)}-1=q_{2}^{(2)}-1=p_{1}^{(22)}=p_{2}^{(22)}=0\right\},$
$P_{2}^{(22)}=\left\{\left(q_{1}^{(2)},q_{2}^{(2)},p_{1}^{(22)},p_{2}^{(22)},t\right)\bigm{|}q_{1}^{(2)}+1=q_{2}^{(2)}-1=p_{1}^{(22)}=p_{2}^{(22)}=0\right\},$
$P_{3}^{(22)}=\left\{\left(q_{1}^{(2)},q_{2}^{(2)},p_{1}^{(22)},p_{2}^{(22)},t\right)\bigm{|}q_{1}^{(2)}-1=q_{2}^{(2)}=p_{1}^{(22)}=p_{2}^{(22)}=0\right\},$
$P_{4}^{(22)}=\left\{\left(q_{1}^{(2)},q_{2}^{(2)},p_{1}^{(22)},p_{2}^{(22)},t\right)\bigm{|}q_{1}^{(2)}-1=q_{2}^{(2)}=(t-1)p_{1}^{(22)}-2=p_{2}^{(22)}=0\right\},$
$P_{5}^{(22)}=\left\{\left(q_{1}^{(2)},q_{2}^{(2)},p_{1}^{(22)},p_{2}^{(22)},t\right)\bigm{|}q_{1}^{(2)}=tq_{2}^{(2)}-1=2tp_{1}^{(22)}+t-1=p_{2}^{(22)}=0\right\},$
$P_{6}^{(22)}=\left\{\left(q_{1}^{(2)},q_{2}^{(2)},p_{1}^{(22)},p_{2}^{(22)},t\right)\bigm{|}q_{1}^{(2)}-1=tq_{2}^{(2)}-1=tp_{1}^{(22)}+1=p_{2}^{(22)}=0\right\},$
$P_{7}^{(22)}=\left\{\left(q_{1}^{(2)},q_{2}^{(2)},p_{1}^{(22)},p_{2}^{(22)},t\right)\bigm{|}q_{1}^{(2)}+1=tq_{2}^{(2)}+1=tp_{1}^{(22)}+1=p_{2}^{(22)}=0\right\}$
on $W_{22}\times B$.
This lemma will be shown in Appendix A. In the following, we use a notation of the accessible singularities
$C_{1}^{(01)}=\left\{q_{1}-q_{2}=p_{1}^{(01)}=p_{2}^{(01)}+1=0\right\}$
for the sake of simplicity.
Although the number of the accessible singularities $C_{j}^{(kl)}$ and $P_{j}^{(kl)}$ are 18 and 42 respectively, we can decrease the number by using relations (2.3) and (2.4).
Lemma 3.2.
We have 6 equalities between 2 accessible singularities as
$C_{1}^{(01)}=C_{1}^{(02)},\quad C_{1}^{(11)}=C_{1}^{(12)},\quad C_{1}^{(21)}=C_{1}^{(22)},\quad C_{2}^{(01)}=C_{2}^{(11)},\quad C_{2}^{(02)}=C_{2}^{(22)},\quad C_{2}^{(12)}=C_{2}^{(21)}.$
Proof.
We first prove the equality $C_{1}^{(01)}=C_{1}^{(02)}$. Relation (2.4) implies
$\left(p_{1}^{(01)},p_{2}^{(01)}\right)=\left(\frac{p_{2}^{(02)}}{p_{1}^{(02)}},\frac{1}{p_{1}^{(02)}}\right),$
from which we obtain that two conditions $p_{1}^{(01)}=p_{2}^{(01)}+1=0$ of $C_{1}^{(01)}$ and $p_{1}^{(02)}+1=p_{2}^{(02)}=0$ of $C_{1}^{(02)}$ are equivalent. The equalities $C_{1}^{(11)}=C_{1}^{(12)}$ and $C_{1}^{(21)}=C_{1}^{(22)}$ can be shown in a similar manner.
We next prove the equality $C_{2}^{(01)}=C_{2}^{(11)}$. Relations (2.3) and (2.4) imply
$\begin{split}\left(q_{1},q_{2},p_{1}^{(01)},p_{2}^{(01)}\right)=\left(\frac{1}{q_{1}^{(1)}},\frac{q_{2}^{(1)}}{q_{1}^{(1)}},-\frac{p_{1}^{(11)}}{q_{1}^{(1)}\left(q_{1}^{(1)}+q_{2}^{(1)}p_{2}^{(11)}+\eta p_{1}^{(11)}\right)},-\frac{p_{2}^{(11)}}{q_{1}^{(1)}+q_{2}^{(1)}p_{2}^{(11)}+\eta p_{1}^{(11)}}\right),\end{split}$ (3.1)
from which we obtain that two conditions $q_{1}-1=p_{1}^{(01)}=p_{2}^{(01)}=0$ of $C_{2}^{(01)}$ and $q_{1}^{(1)}-1=p_{1}^{(11)}=p_{2}^{(11)}=0$ of $C_{2}^{(11)}$ are equivalent. The equalities $C_{2}^{(02)}=C_{2}^{(22)}$ and $C_{2}^{(12)}=C_{2}^{(21)}$ can be shown in a similar manner. ∎
Lemma 3.3.
We have 30 equalities between 2 or 4 accessible singularities as
$P_{3}^{(01)}=P_{3}^{(21)},\quad P_{3}^{(02)}=P_{3}^{(12)},\quad P_{3}^{(11)}=P_{3}^{(22)},$
$P_{1}^{(01)}=P_{1}^{(11)}=P_{6}^{(21)}=P_{6}^{(22)},\quad P_{2}^{(01)}=P_{2}^{(11)}=P_{7}^{(21)}=P_{7}^{(22)},\quad P_{1}^{(02)}=P_{6}^{(11)}=P_{6}^{(12)}=P_{1}^{(22)},$
$P_{2}^{(02)}=P_{7}^{(11)}=P_{7}^{(12)}=P_{2}^{(22)},\quad P_{4}^{(01)}=P_{4}^{(02)}=P_{5}^{(21)}=P_{5}^{(22)},\quad P_{5}^{(01)}=P_{5}^{(02)}=P_{4}^{(11)}=P_{4}^{(12)},$
$P_{6}^{(01)}=P_{6}^{(02)}=P_{1}^{(12)}=P_{1}^{(21)},\quad P_{7}^{(01)}=P_{7}^{(02)}=P_{2}^{(12)}=P_{2}^{(21)},\quad P_{5}^{(11)}=P_{5}^{(12)}=P_{4}^{(21)}=P_{4}^{(22)}.$
Proof.
We prove only the equality $P_{1}^{(01)}=P_{1}^{(11)}=P_{6}^{(21)}=P_{6}^{(22)}$ here. We first obtain from relation (3.1) that two conditions $q_{1}-t=q_{2}-t=p_{1}^{(01)}=p_{2}^{(01)}=0$ of $P_{1}^{(01)}$ and $tq_{1}^{(1)}-1=q_{2}^{(1)}-1=p_{1}^{(11)}=p_{2}^{(11)}=0$ of $P_{1}^{(11)}$ are equivalent. Next, relations (2.3) and (2.4) imply
$\left(q_{1},q_{2},p_{1}^{(01)},p_{2}^{(01)}\right)=\left(\frac{q_{1}^{(2)}}{q_{2}^{(2)}},\frac{1}{q_{2}^{(2)}},\frac{p_{1}^{(21)}}{q_{2}^{(2)}},-\left(q_{1}^{(2)}+q_{2}^{(2)}p_{2}^{(21)}+\eta p_{1}^{(21)}\right)\right),$
from which we obtain that the condition $q_{1}^{(2)}-1=tq_{2}^{(2)}-1=p_{1}^{(21)}=p_{2}^{(21)}+t=0$ of $P_{6}^{(21)}$ coincides with that of $P_{1}^{(01)}$. In the last, relation (2.4) implies
$\left(p_{1}^{(21)},p_{2}^{(21)}\right)=\left(\frac{p_{2}^{(22)}}{p_{1}^{(22)}},\frac{1}{p_{1}^{(22)}}\right),$
from which we obtain that the condition $q_{1}^{(2)}-1=tq_{2}^{(2)}-1=tp_{1}^{(22)}+1=p_{2}^{(22)}=0$ of $P_{6}^{(22)}$ coincides with that of $P_{6}^{(21)}$. Hence we obtain the equality
$P_{1}^{(01)}=P_{1}^{(11)}=P_{6}^{(21)}=P_{6}^{(22)}.$
The other equalities can be shown in a similar manner. ∎
We have reduced the accessible singularities of the system $\mathcal{H}$ to
$\begin{split}&C_{1}^{(01)},\ C_{2}^{(01)},\ C_{3}^{(01)},\ C_{2}^{(02)},\ C_{3}^{(02)},\ C_{3}^{(11)},\ C_{1}^{(12)},\ C_{2}^{(12)},\ C_{3}^{(12)},\ C_{3}^{(21)},\ C_{1}^{(22)},\ C_{3}^{(22)},\\ &P_{3}^{(01)},\ P_{4}^{(01)},\ P_{6}^{(01)},\ P_{7}^{(01)},\ P_{3}^{(12)},\ P_{4}^{(12)},\ P_{6}^{(12)},\ P_{7}^{(12)},\ P_{3}^{(22)},\ P_{4}^{(22)},\ P_{6}^{(22)},\ P_{7}^{(22)}\end{split}$ (3.2)
thanks to Lemma 3.2 and Lemma 3.3. Let $C$ be a union of the accessible singularities defined by
$C=C_{1}^{(01)}\cup C_{2}^{(01)}\cup C_{3}^{(01)}\cup C_{2}^{(02)}\cup C_{3}^{(02)}\cup C_{3}^{(11)}\cup C_{1}^{(12)}\cup C_{2}^{(12)}\cup C_{3}^{(12)}\cup C_{3}^{(21)}\cup C_{1}^{(22)}\cup C_{3}^{(22)}.$
Then we obtain the following lemma.
Lemma 3.4.
The union $C$ is described as a disjoint union of 6 connected components
$C=C_{2}^{(01)}\sqcup C_{2}^{(12)}\sqcup C_{2}^{(22)}\sqcup\left(C_{3}^{(01)}\cup C_{3}^{(02)}\right)\sqcup\left(C_{3}^{(11)}\cup C_{3}^{(12)}\right)\sqcup\left(C_{3}^{(21)}\cup C_{3}^{(22)}\right).$
Moreover, $C$ does not contain any accessible singularity $P_{j}^{(kl)}$ appeared in (3.2).
Proof.
We first consider the accessible singularities except for $C_{1}^{(01)},C_{1}^{(12)},C_{1}^{(22)}$. Using the homogeneous coordinates (2.2), we have
$C_{2}^{(01)}$ $=\left\{\left(\xi,\zeta^{(0)},t\right)\bigm{|}\xi_{0}-\xi_{1}=\zeta_{0}^{(0)}=\zeta_{2}^{(0)}=0,\ \xi_{0}\neq 0,\ \zeta_{1}^{(0)}\neq 0\right\}$
$=\left\{\left([\xi_{0}:\xi_{0}:\xi_{2}],[0:1:0],t\right)\bigm{|}\xi_{0}\neq 0\right\},$
$C_{2}^{(12)}$ $=\left\{\left(\xi,\zeta^{(1)},t\right)\bigm{|}\xi_{1}-\xi_{2}=\zeta_{0}^{(1)}=\zeta_{1}^{(1)}=0,\ \xi_{1}\neq 0,\ \zeta_{2}^{(1)}\neq 0\right\}$
$=\left\{\left([\xi_{0}:\xi_{1}:\xi_{1}],[0:0:1],t\right)\bigm{|}\xi_{1}\neq 0\right\},$
$C_{2}^{(22)}$ $=\left\{\left(\xi,\zeta^{(2)},t\right)\bigm{|}t\xi_{0}-\xi_{2}=\zeta_{0}^{(2)}=\zeta_{1}^{(2)}=0,\ \xi_{2}\neq 0,\ \zeta_{2}^{(2)}\neq 0\right\}$
$=\left\{\left(\left[t^{-1}\xi_{2}:\xi_{1}:\xi_{2}\right],[0:0:1],t\right)\bigm{|}\xi_{2}\neq 0\right\},$
$C_{3}^{(01)}$ $=\left\{\left(\xi,\zeta^{(0)},t\right)\bigm{|}\xi_{1}=\xi_{2}=\zeta_{0}^{(0)}=0,\ \xi_{0}\neq 0,\ \zeta_{1}^{(0)}\neq 0\right\}$
$=\left\{\left([1:0:0],\left[0:\zeta_{1}^{(0)}:\zeta_{2}^{(0)}\right],t\right)\bigm{|}\zeta_{1}^{(0)}\neq 0\right\},$
$C_{3}^{(02)}$ $=\left\{\left(\xi,\zeta^{(0)},t\right)\bigm{|}\xi_{1}=\xi_{2}=\zeta_{0}^{(0)}=0,\ \xi_{0}\neq 0,\ \zeta_{2}^{(0)}\neq 0\right\}$
$=\left\{\left([1:0:0],\left[0:\zeta_{1}^{(0)}:\zeta_{2}^{(0)}\right],t\right)\bigm{|}\zeta_{2}^{(0)}\neq 0\right\},$
$C_{3}^{(11)}$ $=\left\{\left(\xi,\zeta^{(1)},t\right)\bigm{|}\xi_{0}=\xi_{2}=\zeta_{0}^{(1)}=0,\ \xi_{1}\neq 0,\ \zeta_{1}^{(1)}\neq 0\right\}$
$=\left\{\left([0:1:0],\left[0:\zeta_{1}^{(1)}:\zeta_{2}^{(1)}\right],t\right)\bigm{|}\zeta_{1}^{(1)}\neq 0\right\},$
$C_{3}^{(12)}$ $=\left\{\left(\xi,\zeta^{(1)},t\right)\bigm{|}\xi_{0}=\xi_{2}=\zeta_{0}^{(1)}=0,\ \xi_{1}\neq 0,\ \zeta_{2}^{(1)}\neq 0\right\}$
$=\left\{\left([0:1:0],\left[0:\zeta_{1}^{(1)}:\zeta_{2}^{(1)}\right],t\right)\bigm{|}\zeta_{2}^{(1)}\neq 0\right\},$
$C_{3}^{(21)}$ $=\left\{\left(\xi,\zeta^{(2)},t\right)\bigm{|}\xi_{0}=\xi_{1}=\zeta_{0}^{(2)}=0,\ \xi_{2}\neq 0,\ \zeta_{1}^{(2)}\neq 0\right\}$
$=\left\{\left([0:0:1],\left[0:\zeta_{1}^{(2)}:\zeta_{2}^{(2)}\right],t\right)\bigm{|}\zeta_{1}^{(2)}\neq 0\right\},$
$C_{3}^{(22)}$ $=\left\{\left(\xi,\zeta^{(2)},t\right)\bigm{|}\xi_{0}=\xi_{1}=\zeta_{0}^{(2)}=0,\ \xi_{2}\neq 0,\ \zeta_{2}^{(2)}\neq 0\right\}$
$=\left\{\left([0:0:1],\left[0:\zeta_{1}^{(2)}:\zeta_{2}^{(2)}\right],t\right)\bigm{|}\zeta_{2}^{(2)}\neq 0\right\}.$
It implies
$\left(C_{3}^{(01)}\cup C_{3}^{(02)}\right)\cap\left(C_{3}^{(11)}\cup C_{3}^{(12)}\right)=\emptyset,$
$\left(C_{3}^{(01)}\cup C_{3}^{(02)}\right)\cap\left(C_{3}^{(21)}\cup C_{3}^{(22)}\right)=\emptyset,$
$\left(C_{3}^{(11)}\cup C_{3}^{(12)}\right)\cap\left(C_{3}^{(21)}\cup C_{3}^{(22)}\right)=\emptyset$
and
$C_{2}^{(01)}\cap\left(C_{3}^{(01)}\cup C_{3}^{(02)}\right)=\emptyset,\quad C_{2}^{(01)}\cap\left(C_{3}^{(11)}\cup C_{3}^{(12)}\right)=\emptyset,\quad C_{2}^{(01)}\cap\left(C_{3}^{(21)}\cup C_{3}^{(22)}\right)=\emptyset,$
$C_{2}^{(12)}\cap\left(C_{3}^{(01)}\cup C_{3}^{(02)}\right)=\emptyset,\quad C_{2}^{(12)}\cap\left(C_{3}^{(11)}\cup C_{3}^{(12)}\right)=\emptyset,\quad C_{2}^{(12)}\cap\left(C_{3}^{(21)}\cup C_{3}^{(22)}\right)=\emptyset,$
$C_{2}^{(22)}\cap\left(C_{3}^{(01)}\cup C_{3}^{(02)}\right)=\emptyset,\quad C_{2}^{(22)}\cap\left(C_{3}^{(11)}\cup C_{3}^{(12)}\right)=\emptyset,\quad C_{2}^{(22)}\cap\left(C_{3}^{(21)}\cup C_{3}^{(22)}\right)=\emptyset.$
Note that
$C_{3}^{(01)}\cup C_{3}^{(02)}\simeq\mathbb{P}^{1},\quad C_{3}^{(11)}\cup C_{3}^{(12)}\simeq\mathbb{P}^{1},\quad C_{3}^{(21)}\cup C_{3}^{(22)}\simeq\mathbb{P}^{1}.$
Moreover, the accessible singularity $C_{2}^{(01)}$ is transformed to
$\left\{\left(\xi,\zeta^{(1)},t\right)\bigm{|}\xi_{0}-\xi_{1}=\zeta_{0}^{(1)}=\zeta_{2}^{(1)}=0,\ \xi_{1}\neq 0,\ \zeta_{1}^{(1)}\neq 0\right\}=\left\{\left([\xi_{1}:\xi_{1}:\xi_{2}],[0:1:0],t\right)\bigm{|}\xi_{1}\neq 0\right\}$
via relation (2.1), from which we obtain $C_{2}^{(01)}\cap C_{2}^{(12)}=\emptyset$. We also obtain
$C_{2}^{(01)}\cap C_{2}^{(22)}=\emptyset,\quad C_{2}^{(12)}\cap C_{2}^{(22)}=\emptyset$
in a similar manner.
We next consider the rest three accessible singularities. Using the homogeneous coordinates (2.2), we have
$C_{1}^{(01)}$ $=\left\{\left(\xi,\zeta^{(0)},t\right)\bigm{|}\xi_{1}-\xi_{2}=\zeta_{0}^{(0)}=\zeta_{1}^{(0)}+\zeta_{2}^{(0)}=0,\ \xi_{0}\neq 0,\ \zeta_{1}^{(0)}\neq 0\right\}$
$=\left\{\left([\xi_{0}:\xi_{1}:\xi_{1}],[0:1:-1],t\right)\bigm{|}\xi_{0}\neq 0\right\}.$
Under the condition $\xi_{1}\neq 0$, via relation (2.1), the accessible singularity $C_{1}^{(01)}$ is transformed to
$\left\{\left(\xi,\zeta^{(1)},t\right)\bigm{|}\xi_{1}-\xi_{2}=\zeta_{0}^{(1)}=\zeta_{1}^{(1)}=0,\ \xi_{0}\neq 0,\ \xi_{1}\neq 0,\ \zeta_{2}^{(1)}\neq 0\right\}$
$=\left\{\left([\xi_{0}:\xi_{1}:\xi_{1}],[0:0:1],t\right)\bigm{|}\xi_{0}\neq 0,\ \xi_{1}\neq 0\right\},$
which is contained in $C_{2}^{(12)}$. Under the condition $\xi_{1}=0$, $C_{1}^{(01)}$ is equivalent to
$\left\{\left(\xi,\zeta^{(0)},t\right)\bigm{|}\xi_{1}=\xi_{2}=\zeta_{0}^{(0)}=\zeta_{1}^{(0)}+\zeta_{2}^{(0)}=0,\ \xi_{0}\neq 0,\ \zeta_{1}^{(0)}\neq 0\right\}=\left\{\left([1:0:0],[0:1:-1],t\right)\right\},$
which is contained in $C_{3}^{(01)}\cup C_{3}^{(02)}$. Hence we obtain
$C_{1}^{(01)}\subset C_{2}^{(12)}\sqcup\left(C_{3}^{(01)}\cup C_{3}^{(02)}\right).$
We also obtain
$C_{1}^{(12)}\subset C_{2}^{(22)}\sqcup\left(C_{3}^{(11)}\cup C_{3}^{(12)}\right),\quad C_{1}^{(22)}\subset C_{2}^{(01)}\sqcup\left(C_{3}^{(21)}\cup C_{3}^{(22)}\right)$
in a similar manner.
In the last, we consider the accessible singularities $P_{j}^{(kl)}$. Using the homogeneous coordinates (2.2), we have
$P_{3}^{(01)}$ $=\left\{\left(\xi,\zeta^{(0)},t\right)\bigm{|}\xi_{1}=\xi_{2}-t\xi_{0}=\zeta_{0}^{(0)}=\zeta_{2}^{(0)}=0,\ \xi_{0}\neq 0,\ \zeta_{1}^{(0)}\neq 0\right\}$
$=\left\{\left([1:0:t],[0:1:0],t\right)\right\}.$
Recall that $P_{3}^{(01)}$ is equivalent to
$P_{3}^{(21)}$ $=\left\{\left(\xi,\zeta^{(2)},t\right)\bigm{|}\xi_{1}=\xi_{2}-t\xi_{0}=\zeta_{0}^{(2)}=\zeta_{2}^{(2)}=0,\ \xi_{2}\neq 0,\ \zeta_{1}^{(2)}\neq 0\right\}$
$=\left\{\left([1:0:t],[0:1:0],t\right)\right\}.$
We can see that $P_{3}^{(01)}=P_{3}^{(21)}$ is not contained in $C$. We can prove for the other accessible singularities in a similar manner. ∎
Lemma 3.2, Lemma 3.3 and Lemma 3.4 imply the following proposition.
Proposition 3.5.
All of the accessible singularities of the system $\mathcal{H}$ is reduced to
$C_{2}^{(01)},\ C_{3}^{(01)},\ C_{3}^{(02)},\ C_{3}^{(11)},\ C_{2}^{(12)},\ C_{3}^{(12)},\ C_{3}^{(21)},\ C_{2}^{(22)},\ C_{3}^{(22)}$ (3.3)
and
$P_{3}^{(01)},\ P_{4}^{(01)},\ P_{6}^{(01)},\ P_{7}^{(01)},\ P_{3}^{(12)},\ P_{4}^{(12)},\ P_{6}^{(12)},\ P_{7}^{(12)},\ P_{3}^{(22)},\ P_{4}^{(22)},\ P_{6}^{(22)},\ P_{7}^{(22)}.$ (3.4)
4 Resolving the accessible singularities $C_{j}^{(kl)}$
In the previous section, we have determined the accessible singularities of the system $\mathcal{H}$. As will be seen below, there are infinitely many solutions passing through each accessible singularity. In this section, we resolve the accessible singularities (3.3) by introducing new coordinates which parametrize these solutions. We give the result of this section in advance.
Proposition 4.1.
Coordinates resolving the accessible singularities (3.3) are given as follows.
$C_{2}^{(01)}:\left(x_{2}^{[01]},y_{2}^{[01]},z_{2}^{[01]},w_{2}^{[01]}\right)=\left(-\{(q_{1}-1)p_{1}-\alpha_{0}\}p_{1},q_{2},\dfrac{1}{p_{1}},p_{2}\right).$
$C_{3}^{(01)}:\left(x_{3}^{[01]},y_{3}^{[01]},z_{3}^{[01]},w_{3}^{[01]}\right)=\left(-(q_{1}p_{1}+q_{2}p_{2}-\alpha_{5}+\eta)p_{1},q_{2}p_{1},\dfrac{1}{p_{1}},\dfrac{p_{2}}{p_{1}}\right).$
$C_{3}^{(02)}:\left(x_{3}^{[02]},y_{3}^{[02]},z_{3}^{[02]},w_{3}^{[02]}\right)=\left(q_{1}p_{2},-(q_{1}p_{1}+q_{2}p_{2}-\alpha_{5}+\eta)p_{2},\dfrac{p_{1}}{p_{2}},\dfrac{1}{p_{2}}\right).$
$C_{3}^{(11)}:\left(x_{3}^{[11]},y_{3}^{[11]},z_{3}^{[11]},w_{3}^{[11]}\right)=\left(-\left(q_{1}^{(1)}p_{1}^{(1)}+q_{2}^{(1)}p_{2}^{(1)}-\alpha_{1}+\eta\right)p_{1}^{(1)},q_{2}^{(1)}p_{1}^{(1)},\dfrac{1}{p_{1}^{(1)}},\dfrac{p_{2}^{(1)}}{p_{1}^{(1)}}\right).$
$C_{2}^{(12)}:\left(x_{2}^{[12]},y_{2}^{[12]},z_{2}^{[12]},w_{2}^{[12]}\right)=\left(q_{1}^{(1)},-\left\{\left(q_{2}^{(1)}-1\right)p_{2}^{(1)}-\alpha_{2}\right\}p_{2}^{(1)},p_{1}^{(1)},\dfrac{1}{p_{2}^{(1)}}\right).$
$C_{3}^{(12)}:\left(x_{3}^{[12]},y_{3}^{[12]},z_{3}^{[12]},w_{3}^{[12]}\right)=\left(q_{1}^{(1)}p_{2}^{(1)},-\left(q_{1}^{(1)}p_{1}^{(1)}+q_{2}^{(1)}p_{2}^{(1)}-\alpha_{1}+\eta\right)p_{2}^{(1)},\dfrac{p_{1}^{(1)}}{p_{2}^{(1)}},\dfrac{1}{p_{2}^{(1)}}\right).$
$C_{3}^{(21)}:\left(x_{3}^{[21]},y_{3}^{[21]},z_{3}^{[21]},w_{3}^{[21]}\right)=\left(-\left(q_{1}^{(2)}p_{1}^{(2)}+q_{2}^{(2)}p_{2}^{(2)}-\alpha_{3}+\eta\right)p_{1}^{(2)},q_{2}^{(2)}p_{1}^{(2)},\dfrac{1}{p_{1}^{(2)}},\dfrac{p_{2}^{(2)}}{p_{1}^{(2)}}\right).$
$C_{2}^{(22)}:\left(x_{2}^{[22]},y_{2}^{[22]},z_{2}^{[22]},w_{2}^{[22]}\right)=\left(q_{1}^{(2)},-\left\{\left(q_{2}^{(2)}-\frac{1}{t}\right)p_{2}^{(2)}-\alpha_{4}\right\}p_{2}^{(2)},p_{1}^{(2)},\frac{1}{p_{2}^{(2)}}\right).$
$C_{3}^{(22)}:\left(x_{3}^{[22]},y_{3}^{[22]},z_{3}^{[22]},w_{3}^{[22]}\right)=\left(q_{1}^{(2)}p_{2}^{(2)},-\left(q_{1}^{(2)}p_{1}^{(2)}+q_{2}^{(2)}p_{2}^{(2)}-\alpha_{3}+\eta\right)p_{2}^{(2)},\dfrac{p_{1}^{(2)}}{p_{2}^{(2)}},\dfrac{1}{p_{2}^{(2)}}\right).$
Namely, the system on each coordinate $\left(x_{j}^{[kl]},y_{j}^{[kl]},z_{j}^{[kl]},w_{j}^{[kl]}\right)$ is described as a polynomial Hamiltonian system.
In this proposition, we denote the coordinate resolving $C_{j}^{(kl)}$ by $\left(x_{j}^{[kl]},y_{j}^{[kl]},z_{j}^{[kl]},w_{j}^{[kl]}\right)$. This proposition is shown in two ways. One is the blowing up method and another is that in terms of formal power series. Note that the latter method will be used in the next section to resolve $P_{j}^{(kl)}$. Because we can prove in the same way for all of the accessible singularities, we consider only $C_{2}^{(01)}=\left\{q_{1}-1=p_{1}^{(01)}=p_{2}^{(01)}=0\right\}$ here. The system $\mathcal{H}$ on $W_{01}\times B$ is described as
$\begin{split}t(t-1)\frac{dq_{1}}{dt}&=(\alpha_{1}+\eta)q_{1}^{2}+\alpha_{3}q_{1}q_{2}+\{(\alpha_{0}+\alpha_{5}-\eta)t-(\alpha_{0}+\alpha_{1}+\eta)\}q_{1}-\alpha_{3}q_{2}-(\alpha_{5}-\eta)t\\ &\quad+\frac{(q_{1}-1)(q_{1}+q_{2})(q_{2}-t)p_{2}^{(01)}+2q_{1}(q_{1}-1)(q_{1}-t)}{p_{1}^{(01)}},\\ t(t-1)\frac{dq_{2}}{dt}&=\alpha_{1}q_{1}q_{2}+(\alpha_{3}+\eta)q_{2}^{2}-\alpha_{1}tq_{1}-\{(\alpha_{3}+\alpha_{4}+\eta-1)t-(\alpha_{4}+\alpha_{5}-\eta-1)\}q_{2}-(\alpha_{5}-\eta)t\\ &\quad+\frac{2q_{2}(q_{2}-1)(q_{2}-t)p_{2}^{(01)}+(q_{1}-1)(q_{1}+q_{2})(q_{2}-t)}{p_{1}^{(01)}},\\ t(t-1)\frac{dp_{1}^{(01)}}{dt}&=\alpha_{1}\eta\left(p_{1}^{(01)}\right)^{2}+\{2(\alpha_{1}+\eta)q_{1}+(\alpha_{1}+\alpha_{3})q_{2}+(\alpha_{0}-\alpha_{1}+\alpha_{5}-\eta)t-(\alpha_{0}+\alpha_{1}+\eta)\}p_{1}^{(01)}\\ &\quad+(2q_{1}+q_{2}-1)(q_{2}-t)p_{2}^{(01)}+3q_{1}^{2}-2(t+1)q_{1}+t,\\ t(t-1)\frac{dp_{2}^{(01)}}{dt}&=\eta\left(\alpha_{1}p_{2}^{(01)}-\alpha_{3}\right)p_{1}^{(01)}+\alpha_{1}(q_{2}-t)\left(p_{2}^{(01)}\right)^{2}\\ &\quad+\{(\alpha_{1}+2\eta)q_{1}-(\alpha_{3}+2\eta)q_{2}-\alpha_{1}t+\alpha_{2}+\alpha_{3}\}p_{2}^{(01)}-\alpha_{3}(q_{1}-1)\\ &\quad+\frac{\left\{2q_{1}q_{2}-2q_{2}^{2}-2tq_{1}+(t+1)q_{2}\right\}\left(p_{2}^{(01)}\right)^{2}+\left\{2q_{1}q_{2}+2q_{2}^{2}-(t+1)q_{1}+2q_{2}\right\}p_{2}^{(01)}}{p_{1}^{(01)}}.\end{split}$ (4.1)
4.1 Blowing up
Let $\left(u_{1},u_{2},v_{1},v_{2}\right)$, $\left(u^{\prime}_{1},u^{\prime}_{2},v^{\prime}_{1},v^{\prime}_{2}\right)$ and $\left(u^{\prime\prime}_{1},u^{\prime\prime}_{2},v^{\prime\prime}_{1},v^{\prime\prime}_{2}\right)$ be affine coordinates of $\mathbb{C}^{4}$ defined by
$\begin{split}&q_{1}-1=u_{1}v_{1},\quad q_{2}=u_{2},\quad p_{1}^{(01)}=v_{1},\quad p_{2}^{(01)}=v_{1}v_{2},\\ &q_{1}-1=u^{\prime}_{1},\quad q_{2}=u^{\prime}_{2},\quad p_{1}^{(01)}=u^{\prime}_{1}v^{\prime}_{1},\quad p_{2}^{(01)}=u^{\prime}_{1}v^{\prime}_{2},\\ &q_{1}-1=u^{\prime\prime}_{1}v^{\prime\prime}_{2},\quad q_{2}=u^{\prime\prime}_{2},\quad p_{1}^{(01)}=v^{\prime\prime}_{1}v^{\prime\prime}_{2},\quad p_{2}^{(01)}=v^{\prime\prime}_{2}.\end{split}$ (4.2)
On these coordinates, the accessible singularity $C_{2}^{(01)}$ corresponds to
$E_{1}=\left\{v_{1}=0\right\},\quad E^{\prime}_{1}=\left\{u^{\prime}_{1}=0\right\},\quad E^{\prime\prime}_{1}=\left\{v^{\prime\prime}_{2}=0\right\}.$
The union $E_{1}\cup E^{\prime}_{1}\cup E^{\prime\prime}_{1}$ is called the exceptional divisor and $E_{1}\cup E^{\prime}_{1}\cup E^{\prime\prime}_{1}\simeq\mathbb{C}^{1}\times\mathbb{P}^{2}$. System (4.1) on the coordinate $(u_{1},u_{2},v_{1},v_{2})$ is described as
$\frac{du_{1}}{dt}=\frac{\alpha_{0}-u_{1}}{tv_{1}}+U_{1},\quad\frac{du_{2}}{dt}=U_{2},\quad\frac{dv_{1}}{dt}=V_{1},\quad\frac{dv_{2}}{dt}=V_{2},$ (4.3)
where $U_{1},U_{2},V_{1},V_{2}\in\mathbb{C}(t)[u_{1},u_{2},v_{1},v_{2}]$ and its accessible singularity is given by
$D_{1}=\left\{\alpha_{0}-u_{1}=v_{1}=0\right\}\subset E_{1}.$
System (4.1) on the coordinate $(u^{\prime}_{1},u^{\prime}_{2},v^{\prime}_{1},v^{\prime}_{2})$ is described as
$\frac{du^{\prime}_{1}}{dt}=\dfrac{U^{\prime}_{1}}{v^{\prime}_{1}},\quad\frac{du^{\prime}_{2}}{dt}=\dfrac{U^{\prime}_{2}}{v^{\prime}_{1}},\quad\frac{dv^{\prime}_{1}}{dt}=\dfrac{1-\alpha_{0}v^{\prime}_{1}}{tu^{\prime}_{1}}+V^{\prime}_{1},\quad\frac{dv^{\prime}_{2}}{dt}=\dfrac{(t-1)\left(1-\alpha_{0}v^{\prime}_{1}\right)v^{\prime}_{2}+u^{\prime}_{1}V^{\prime}_{2}}{t(t-1)u^{\prime}_{1}v^{\prime}_{1}},$
where $U^{\prime}_{1},U^{\prime}_{2},V^{\prime}_{1},V^{\prime}_{2}\in\mathbb{C}(t)\left[u^{\prime}_{1},u^{\prime}_{2},v^{\prime}_{1},v^{\prime}_{2}\right]$ and its accessible singularity is given by
$D^{\prime}_{1}=\left\{1-\alpha_{0}v^{\prime}_{1}=u^{\prime}_{1}=0\right\}\subset E^{\prime}_{1}.$
System (4.1) on the coordinate $(u^{\prime\prime}_{1},u^{\prime\prime}_{2},v^{\prime\prime}_{1},v^{\prime\prime}_{2})$ is described as
$\frac{du^{\prime\prime}_{1}}{dt}=\dfrac{\left(t-1\right)\left(\alpha_{0}v^{\prime\prime}_{1}-u^{\prime\prime}_{1}\right)+v^{\prime\prime}_{2}U^{\prime\prime}_{1}}{t\left(t-1\right)v^{\prime\prime}_{1}v^{\prime\prime}_{2}},\quad\frac{du^{\prime\prime}_{2}}{dt}=\dfrac{U^{\prime\prime}_{2}}{v^{\prime\prime}_{1}},\quad\frac{dv^{\prime\prime}_{1}}{dt}=V^{\prime\prime}_{1},\quad\frac{dv^{\prime\prime}_{2}}{dt}=\dfrac{1-t+v^{\prime\prime}_{2}V^{\prime\prime}_{2}}{t\left(t-1\right)v^{\prime\prime}_{1}},$
where $U^{\prime\prime}_{1},U^{\prime\prime}_{2},V^{\prime\prime}_{1},V^{\prime\prime}_{2}\in\mathbb{C}(t)\left[u^{\prime\prime}_{1},u^{\prime\prime}_{2},v^{\prime\prime}_{1},v^{\prime\prime}_{2}\right]$ and its accessible singularity is given by
$D^{\prime\prime}_{1}=\left\{\alpha_{0}v^{\prime\prime}_{1}-u^{\prime\prime}_{1}=v^{\prime\prime}_{2}=0,\ v^{\prime\prime}_{1}\neq 0\right\}\subset E^{\prime\prime}_{1}.$
Relation (4.2) implies
$\left(u^{\prime}_{1},u^{\prime}_{2},v^{\prime}_{1},v^{\prime}_{2}\right)=\left(u_{1}v_{1},u_{2},\frac{1}{u_{1}},\frac{v_{2}}{u_{1}}\right),\quad\left(u^{\prime\prime}_{1},u^{\prime\prime}_{2},v^{\prime\prime}_{1},v^{\prime\prime}_{2}\right)=\left(\frac{u_{1}}{v_{2}},u_{2},\frac{1}{v_{2}},v_{1}v_{2}\right),$
from which we have $D^{\prime}_{1}=D_{1}$ and $D^{\prime\prime}_{1}\subset D_{1}$. Note that $D_{1}\simeq\mathbb{C}^{2}$. Hence it is enough to consider system (4.3) and its accessible singularity $D_{1}$.
Let $\left(x_{2}^{[01]},y_{2}^{[01]},z_{2}^{[01]},w_{2}^{[01]}\right)$ and $(x^{\prime}_{2},y^{\prime}_{2},z^{\prime}_{2},w^{\prime}_{2})$ be affine coordinates of $\mathbb{C}^{4}$ defined by
$\begin{split}&\alpha_{0}-u_{1}=x_{2}^{[01]}z_{2}^{[01]},\quad u_{2}=y_{2}^{[01]},\quad v_{1}=z_{2}^{[01]},\quad v_{2}=w_{2}^{[01]},\\ &\alpha_{0}-u_{1}=x^{\prime}_{2},\quad u_{2}=y^{\prime}_{2},\quad v_{1}=x^{\prime}_{2}z^{\prime}_{2},\quad v_{2}=w^{\prime}_{2}.\end{split}$ (4.4)
On these coordinates, the accessible singularity $D_{1}$ corresponds to
$E_{2}=\left\{z_{2}^{[01]}=0\right\},\quad E^{\prime}_{2}=\left\{x^{\prime}_{2}=0\right\}.$
Note that $E_{2}\cup E^{\prime}_{2}\simeq\mathbb{C}^{2}\times\mathbb{P}^{1}$. System (4.3) on the coordinate $\left(x_{2}^{[01]},y_{2}^{[01]},z_{2}^{[01]},w_{2}^{[01]}\right)$ is described as
$\frac{dx_{2}^{[01]}}{dt}=X_{2},\quad\frac{dy_{2}^{[01]}}{dt}=Y_{2},\quad\frac{dz_{2}^{[01]}}{dt}=Z_{2},\quad\frac{dw_{2}^{[01]}}{dt}=W_{2},$ (4.5)
where $X_{2},Y_{2},Z_{2},W_{2}\in\mathbb{C}(t)\left[x_{2}^{[01]},y_{2}^{[01]},z_{2}^{[01]},w_{2}^{[01]}\right]$ and it does not have any singularity. System (4.3) on the coordinate $\left(x^{\prime}_{2},y^{\prime}_{2},z^{\prime}_{2},w^{\prime}_{2}\right)$ is described as
$\frac{dx^{\prime}_{2}}{dt}=\dfrac{1-t+z^{\prime}_{2}X^{\prime}_{2}}{t(t-1)z^{\prime}_{2}},\quad\frac{dy^{\prime}_{2}}{dt}=Y^{\prime}_{2},\quad\frac{dz^{\prime}_{2}}{dt}=Z^{\prime}_{2},\quad\frac{dw^{\prime}_{2}}{dt}=W^{\prime}_{2},$
where $X^{\prime}_{2},Y^{\prime}_{2},Z^{\prime}_{2},W^{\prime}_{2}\in\mathbb{C}(t)\left[x^{\prime}_{2},y^{\prime}_{2},z^{\prime}_{2},w^{\prime}_{2}\right]$ and it does not have any accessible singularity. Note that the condition $z^{\prime}_{2}=0$ corresponds to $x^{[01]}_{2}$$\infty$ and the solution satisfying $z^{\prime}_{2}=0$ does not pass through $\Sigma_{\eta}^{0}$. Hence it is enough to consider system (4.5).
The relation between the coordinates $\left(q_{1},q_{2},p_{1}^{(01)},p_{2}^{(01)}\right)$ and $\left(x_{2}^{[01]},y_{2}^{[01]},z_{2}^{[01]},w_{2}^{[01]}\right)$ is given by
$q_{1}=1+\left(\alpha_{0}-x_{2}^{[01]}z_{2}^{[01]}\right)z_{2}^{[01]},\quad q_{2}=y_{2}^{[01]},\quad p_{1}^{(01)}=z_{2}^{[01]},\quad p_{2}^{(01)}=z_{2}^{[01]}w_{2}^{[01]}.$ (4.6)
Then we see that the accessible singularity $C_{2}^{(01)}$ corresponds to $\left\{z_{2}^{[02]}=0\right\}$. Hence all of the solutions to system (4.1) passing through $C_{2}^{(01)}$ can be parametrized in terms of an initial condition
$\left(x_{2}^{[01]},y_{2}^{[01]},z_{2}^{[01]},w_{2}^{[01]}\right)=(a_{0},b_{0},0,d_{0}),$ (4.7)
where $a_{0},b_{0},d_{0}\in\mathbb{C}$, for system (4.5). Combining relations (2.4) and (4.6), we obtain
$x_{2}^{[01]}=-\{(q_{1}-1)p_{1}-\alpha_{0}\}p_{1},\quad y_{2}^{[01]}=q_{2},\quad z_{2}^{[01]}=\dfrac{1}{p_{1}},\quad w_{2}^{[01]}=p_{2}$ (4.8)
and it is a symplectic transformation. Moreover, system (4.5) is described as a polynomial Hamiltonian system. We can show it by a direct calculation.
4.2 Formal power series
We fix a point $t_{0}\in B$ and consider the solution to system (4.1) with the initial condition
$\left(q_{1}(t_{0}),q_{2}(t_{0}),p_{1}^{(01)}(t_{0}),p_{2}^{(01)}(t_{0})\right)=(1,b_{0},0,0),$
where $b_{0}\in\mathbb{C}$. This solution is expressed in terms of formal power series as
$\begin{split}&q_{1}=1+(t-t_{0})^{\lambda_{a}}\sum_{k=1}^{\infty}a_{k}(t-t_{0})^{k},\quad q_{2}=b_{0}+(t-t_{0})^{\lambda_{b}}\sum_{k=1}^{\infty}b_{k}(t-t_{0})^{k},\\ &p_{1}^{(01)}=(t-t_{0})^{\lambda_{c}}\sum_{k=1}^{\infty}c_{k}(t-t_{0})^{k},\quad p_{2}^{(01)}=(t-t_{0})^{\lambda_{d}}\sum_{k=1}^{\infty}d_{k}(t-t_{0})^{k},\end{split}$ (4.9)
where $\lambda_{a},a_{k},\lambda_{b},b_{k},\lambda_{c},c_{k},\lambda_{d},d_{k}\in\mathbb{C}$ and
${\rm{Re}}(\lambda_{a})>-1,\quad a_{1}\neq 0,\quad{\rm{Re}}(\lambda_{b})>-1,\quad b_{1}\neq 0,\quad{\rm{Re}}(\lambda_{c})>-1,\quad c_{1}\neq 0,\quad{\rm{Re}}(\lambda_{d})>-1,\quad d_{1}\neq 0.$
Substituting solution (4.9) to the third equation of system (4.1), we obtain
$t_{0}(t_{0}-1)(\lambda_{c}+1)c_{1}(t-t_{0})^{\lambda_{c}}+O\left((t-t_{0})^{\lambda_{c}+1}\right)=-t_{0}+1+o(1).$
Here we denote the function $f(t)$ satisfying $\lim_{t\to t_{0}}\left|\frac{f(t)}{(t-t_{0})^{k}}\right|<\infty$ and $\lim_{t\to t_{0}}\left|\frac{f(t)}{(t-t_{0})^{k}}\right|=0$ by $O\left((t-t_{0})^{k}\right)$ and $o\left((t-t_{0})^{k}\right)$ respectively. It follows that
$\lambda_{c}=0,\quad c_{1}=-\frac{1}{t_{0}}.$ (4.10)
Similarly we obtain
$\lambda_{a}=0,\quad a_{1}=-\frac{\alpha_{0}}{t_{0}}$ (4.11)
from the first equation of system (4.1),
$\lambda_{b}=0$ (4.12)
from the second equation and
$\lambda_{d}=0$ (4.13)
from the fourth equation. Under conditions (4.10), (4.11), (4.12) and (4.13), we also obtain
$b_{1}=-\frac{2b_{0}(b_{0}-1)(b_{0}-t_{0})}{t_{0}-1}d_{1}+B_{1},$
where $B_{1}\in\mathbb{C}(t_{0})[\alpha_{0},\ldots,\alpha_{5},\eta,b_{0}]$, from the second equation. Then we see that $d_{1}$ turns to be an arbitrary complex number and $b_{1}$ is determined uniquely by $d_{1}$. Moreover, in a similar manner, $c_{2}$ and $d_{2}$ are determined uniquely as elements of $\mathbb{C}(t_{0})[\alpha_{0},\ldots,\alpha_{5},\eta,b_{0},d_{1}]$ and we obtain
$b_{2}=-\frac{(b_{0}+1)(b_{0}-t_{0})}{2(t_{0}-1)}a_{2}+B_{2},$
where $B_{2}\in\mathbb{C}(t_{0})[\alpha_{0},\ldots,\alpha_{5},\eta,b_{0},d_{1}]$. Then we see that $a_{2}$ turns to be an arbitrary complex number and $b_{2}$ is determined uniquely by $a_{2}$.
By using solution (4.9) with (4.10), (4.11), (4.12) and (4.13), we have
$\frac{q_{1}-1}{p_{1}^{(01)}}$ $=\frac{-\alpha_{0}t_{0}^{-1}(t-t_{0})+a_{2}(t-t_{0})^{2}+O\left((t-t_{0})^{3}\right)}{-t_{0}^{-1}(t-t_{0})+c_{2}(t-t_{0})^{2}+O\left((t-t_{0})^{3}\right)}$
$=\frac{-\alpha_{0}t_{0}^{-1}+a_{2}(t-t_{0})+O\left((t-t_{0})^{2}\right)}{-t_{0}^{-1}+c_{2}(t-t_{0})+O\left((t-t_{0})^{2}\right)}$
$=\alpha_{0}+\frac{-t_{0}^{-1}a_{2}+\alpha_{0}t_{0}^{-1}c_{2}}{t_{0}^{-2}}(t-t_{0})+O\left((t-t_{0})^{2}\right)$
$=\alpha_{0}+t_{0}(-a_{2}+\alpha_{0}c_{2})(t-t_{0})+O\left((t-t_{0})^{2}\right),$
from which we obtain
$\frac{\frac{q_{1}-1}{p_{1}^{(01)}}-\alpha_{0}}{p_{1}^{(01)}}$ $=\frac{t_{0}(-a_{2}+\alpha_{0}c_{2})(t-t_{0})+O\left((t-t_{0})^{2}\right)}{-t_{0}^{-1}(t-t_{0})+O\left((t-t_{0})^{2}\right)}$
$=t_{0}^{2}(a_{2}-\alpha_{0}c_{2})+O(t-t_{0}).$
We also have
$\frac{p_{2}^{(01)}}{p_{1}^{(01)}}=-t_{0}d_{1}+O(t-t_{0})$
in a similar manner. Set
$x_{2}^{[01]}=-\frac{\frac{q_{1}-1}{p_{1}^{(01)}}-\alpha_{0}}{p_{1}^{(01)}},\quad y_{2}^{[01]}=q_{2},\quad z_{2}^{[01]}=p_{1}^{(01)},\quad w_{2}^{[01]}=\frac{p_{2}^{(01)}}{p_{1}^{(01)}}.$ (4.14)
Then we obtain
$\left(x_{2}^{[01]}(t_{0}),y_{2}^{[01]}(t_{0}),z_{2}^{[01]}(t_{0}),w_{2}^{[01]}(t_{0})\right)=\left(t_{0}^{2}(a_{2}-\alpha_{0}c_{2}),b_{0},0,-t_{0}d_{1}\right),$
which is essentially equivalent to (4.7). Note that relation (4.14) is equivalent to (4.6).
5 Resolving the accessible singularities $P_{j}^{(kl)}$
It is not suitable to use the blowing up method to resolve the accessible singularities (3.4), because the explicit formula of the system after blowing up becomes more complicated. Therefore we use the method in terms of formal power series and the symmetry of system (1.1) in this section. We give the result of this section in advance.
Proposition 5.1.
Coordinates resolving the accessible singularities (3.4) are given as follows.
$P_{3}^{(01)}:\left(\widetilde{x}_{3}^{[01]},\widetilde{y}_{3}^{[01]},\widetilde{z}_{3}^{[01]},\widetilde{w}_{3}^{[01]}\right)=r_{3}\left(x_{3}^{[01]},y_{3}^{[01]},z_{3}^{[01]},w_{3}^{[01]}\right).$
$P_{4}^{(01)}:\left(\widetilde{x}_{4}^{[01]},\widetilde{y}_{4}^{[01]},\widetilde{z}_{4}^{[01]},\widetilde{w}_{4}^{[01]}\right)=r_{4}r_{3}\left(x_{3}^{[01]},y_{3}^{[01]},z_{3}^{[01]},w_{3}^{[01]}\right).$
$P_{6}^{(01)}:\left(\widetilde{x}_{6}^{[01]},\widetilde{y}_{6}^{[01]},\widetilde{z}_{6}^{[01]},\widetilde{w}_{6}^{[01]}\right)=r_{4}\left(x_{2}^{[01]},y_{2}^{[01]},z_{2}^{[01]},w_{2}^{[01]}\right).$
$P_{7}^{(01)}:\left(\widetilde{x}_{7}^{[01]},\widetilde{y}_{7}^{[01]},\widetilde{z}_{7}^{[01]},\widetilde{w}_{7}^{[01]}\right)=r_{5}r_{4}\left(x_{2}^{[01]},y_{2}^{[01]},z_{2}^{[01]},w_{2}^{[01]}\right).$
$P_{3}^{(12)}:\left(\widetilde{x}_{3}^{[12]},\widetilde{y}_{3}^{[12]},\widetilde{z}_{3}^{[12]},\widetilde{w}_{3}^{[12]}\right)=r_{5}\left(x_{3}^{[12]},y_{3}^{[12]},z_{3}^{[12]},w_{3}^{[12]}\right).$
$P_{4}^{(12)}:\left(\widetilde{x}_{4}^{[12]},\widetilde{y}_{4}^{[12]},\widetilde{z}_{4}^{[12]},\widetilde{w}_{4}^{[12]}\right)=r_{0}r_{5}\left(x_{3}^{[12]},y_{3}^{[12]},z_{3}^{[12]},w_{3}^{[12]}\right).$
$P_{6}^{(12)}:\left(\widetilde{x}_{6}^{[12]},\widetilde{y}_{6}^{[12]},\widetilde{z}_{6}^{[12]},\widetilde{w}_{6}^{[12]}\right)=r_{0}\left(x_{2}^{[12]},y_{2}^{[12]},z_{2}^{[12]},w_{2}^{[12]}\right).$
$P_{7}^{(12)}:\left(\widetilde{x}_{7}^{[12]},\widetilde{y}_{7}^{[12]},\widetilde{z}_{7}^{[12]},\widetilde{w}_{7}^{[12]}\right)=r_{1}r_{0}\left(x_{2}^{[12]},y_{2}^{[12]},z_{2}^{[12]},w_{2}^{[12]}\right).$
$P_{3}^{(22)}:\left(\widetilde{x}_{3}^{[22]},\widetilde{y}_{3}^{[22]},\widetilde{z}_{3}^{[22]},\widetilde{w}_{3}^{[22]}\right)=r_{1}\left(x_{3}^{[22]},y_{3}^{[22]},z_{3}^{[22]},w_{3}^{[22]}\right).$
$P_{4}^{(22)}:\left(\widetilde{x}_{4}^{[22]},\widetilde{y}_{4}^{[22]},\widetilde{z}_{4}^{[22]},\widetilde{w}_{4}^{[22]}\right)=r_{2}r_{1}\left(x_{3}^{[22]},y_{3}^{[22]},z_{3}^{[22]},w_{3}^{[22]}\right).$
$P_{6}^{(22)}:\left(\widetilde{x}_{6}^{[22]},\widetilde{y}_{6}^{[22]},\widetilde{z}_{6}^{[22]},\widetilde{w}_{6}^{[22]}\right)=r_{2}\left(x_{2}^{[22]},y_{2}^{[22]},z_{2}^{[22]},w_{2}^{[22]}\right).$
$P_{7}^{(22)}:\left(\widetilde{x}_{7}^{[22]},\widetilde{y}_{7}^{[22]},\widetilde{z}_{7}^{[22]},\widetilde{w}_{7}^{[22]}\right)=r_{3}r_{2}\left(x_{2}^{[22]},y_{2}^{[22]},z_{2}^{[22]},w_{2}^{[22]}\right).$
Namely, the system on each coordinate $\left(\widetilde{x}_{j}^{[kl]},\widetilde{y}_{j}^{[kl]},\widetilde{z}_{j}^{[kl]},\widetilde{w}_{j}^{[kl]}\right)$ is described as a polynomial Hamiltonian system.
In this proposition, we denote the coordinates resolving $P_{j}^{(kl)}$ by $\left(\widetilde{x}_{j}^{[kl]},\widetilde{y}_{j}^{[kl]},\widetilde{z}_{j}^{[kl]},\widetilde{w}_{j}^{[kl]}\right)$. We also use the notation
$r_{3}\left(x_{3}^{[01]},y_{3}^{[01]},z_{3}^{[01]},w_{3}^{[01]}\right)$ $=\left(r_{3}\left(-(q_{1}p_{1}+q_{2}p_{2}-\alpha_{5}+\eta)p_{1}\right),r_{3}(q_{2}p_{1}),r_{3}\left(\dfrac{1}{p_{1}}\right),r_{3}\left(\dfrac{p_{2}}{p_{1}}\right)\right)$
$=\left(-(q_{1}p_{1}+q_{2}p_{2}-\alpha_{5}+\eta)p_{1},\frac{(q_{2}p_{2}+\alpha_{3})p_{1}}{p_{2}},\frac{1}{p_{1}},\frac{p_{2}}{p_{1}}\right)$
for the sake of simplicity. Recall that the coordinate $\left(x_{3}^{[01]},y_{3}^{[01]},z_{3}^{[01]},w_{3}^{[01]}\right)$ was given in Proposition 4.1 as
$\left(x_{3}^{[01]},y_{3}^{[01]},z_{3}^{[01]},w_{3}^{[01]}\right)=\left(-(q_{1}p_{1}+q_{2}p_{2}-\alpha_{5}+\eta)p_{1},q_{2}p_{1},\dfrac{1}{p_{1}},\dfrac{p_{2}}{p_{1}}\right).$
5.1 Proof of Proposition 5.1
Because we can prove in the same way for all of the accessible singularities, we consider only $P_{3}^{(01)}=\left\{q_{1}=q_{2}-t=p_{1}^{(01)}=p_{2}^{(01)}=0\right\}$ here. We fix a point $t_{0}\in B$ and consider the solution to system (4.1) with the initial condition
$\left(q_{1}(t_{0}),q_{2}(t_{0}),p_{1}^{(01)}(t_{0}),p_{2}^{(01)}(t_{0})\right)=(0,t_{0},0,0).$ (5.1)
This solution is expressed in terms of formal power series as
$\begin{split}&q_{1}=(t-t_{0})^{\lambda_{a}}\sum_{k=1}^{\infty}a_{k}(t-t_{0})^{k},\quad q_{2}=t_{0}+(t-t_{0})^{\lambda_{b}}\sum_{k=1}^{\infty}b_{k}(t-t_{0})^{k},\\ &p_{1}^{(01)}=(t-t_{0})^{\lambda_{c}}\sum_{k=1}^{\infty}c_{k}(t-t_{0})^{k},\quad p_{2}^{(01)}=(t-t_{0})^{\lambda_{d}}\sum_{k=1}^{\infty}d_{k}(t-t_{0})^{k},\end{split}$ (5.2)
where $\lambda_{a},a_{k},\lambda_{b},b_{k},\lambda_{c},c_{k},\lambda_{d},d_{k}\in\mathbb{C}$ and
${\rm{Re}}(\lambda_{a})>-1,\quad a_{1}\neq 0,\quad{\rm{Re}}(\lambda_{b})>-1,\quad b_{1}\neq 0,\quad{\rm{Re}}(\lambda_{c})>-1,\quad c_{1}\neq 0,\quad{\rm{Re}}(\lambda_{d})>-1,\quad d_{1}\neq 0.$
Substituting solution (5.2) to the third equation of system (4.1), we obtain
$t_{0}(t_{0}-1)(\lambda_{c}+1)c_{1}(t-t_{0})^{\lambda_{c}}+O\left((t-t_{0})^{\lambda_{c}+1}\right)=t_{0}+o(1).$
It follows that
$\lambda_{c}=0,\quad c_{1}=\frac{1}{t_{0}-1}.$ (5.3)
Similarly we obtain
$\lambda_{d}=0,\quad d_{1}=\dfrac{-\alpha_{3}}{t_{0}(t_{0}-1)}$ (5.4)
from the fourth equation of system (4.1),
$\lambda_{a}=0,\quad a_{1}=\dfrac{\alpha_{3}+\alpha_{5}-\eta}{t_{0}-1}$ (5.5)
from the first equation and
$\lambda_{b}=0,\quad b_{1}=1-\dfrac{\alpha_{4}}{2}$ (5.6)
from the second equation. Moreover, in a similar manner, $b_{2}$ and $c_{2}$ are determined uniquely as elements of $\mathbb{C}(t_{0})[\alpha_{0},\ldots,\alpha_{5},\eta]$ and $a_{2}$ and $d_{2}$ turn to be arbitrary complex numbers.
We consider the action of the transformation $r_{3}$ on the above solution. Set
$\widetilde{q}_{1}=r_{3}(q_{1}),\quad\widetilde{q}_{2}=r_{3}(q_{2}),\quad\widetilde{p}_{1}=r_{3}(p_{1}),\quad\widetilde{p}_{2}=r_{3}(p_{2}),\quad\widetilde{p}_{1}^{(01)}=r_{3}\left(p_{1}^{(01)}\right),\quad\widetilde{p}_{2}^{(01)}=r_{3}\left(p_{2}^{(01)}\right).$
Then relation (2.4) implies
$\begin{split}&\widetilde{q}_{1}=q_{1},\\ &\widetilde{q}_{2}=q_{2}+\frac{\alpha_{3}}{p_{2}}=q_{2}+\alpha_{3}\frac{p_{1}^{(01)}}{p_{2}^{(01)}},\\ &\widetilde{p}_{1}^{(01)}=r_{3}\left(\frac{1}{p_{1}}\right)=\frac{1}{p_{1}}=p_{1}^{(01)},\\ &\widetilde{p}_{2}^{(01)}=r_{3}\left(\frac{p_{2}}{p_{1}}\right)=\frac{p_{2}}{p_{1}}=p_{2}^{(01)}.\end{split}$ (5.7)
Substituting solution (5.2) with (5.3), (5.4), (5.5) and (5.6) to relation (5.7), we obtain
$\widetilde{q}_{1}=O(t-t_{0}),\quad\widetilde{p}_{1}^{(01)}=O(t-t_{0}),\quad\widetilde{p}_{2}^{(01)}=O(t-t_{0})$
and
$\widetilde{q}_{2}$ $=\left(t_{0}+O(t-t_{0})\right)+\alpha_{3}\frac{c_{1}(t-t_{0})+O\left((t-t_{0})^{2}\right)}{d_{1}(t-t_{0})+O\left((t-t_{0})^{2}\right)}$
$=\left(t_{0}+\alpha_{3}\frac{c_{1}}{d_{1}}\right)+O(t-t_{0})$
$=O(t-t_{0}).$
Hence the solution to system (4.1) with initial condition (5.1) is transformed to that to the system in terms of the coordinate $\left(\widetilde{q}_{1},\widetilde{q}_{2},\widetilde{p}_{1}^{(01)},\widetilde{p}_{2}^{(01)}\right)$ whose initial condition is given by
$\left(\widetilde{q}_{1}(t_{0}),\widetilde{q}_{2}(t_{0}),\widetilde{p}_{1}^{(01)}(t_{0}),\widetilde{p}_{2}^{(01)}(t_{0})\right)=(0,0,0,0).$ (5.8)
Since system (1.1) is invariant under the transformation $r_{3}$, the system in terms of $\left(\widetilde{q}_{1},\widetilde{q}_{2},\widetilde{p}_{1},\widetilde{p}_{2}\right)$ is the same as system (1.1) except changing the parameters. Hence the system in terms of $\left(\widetilde{q}_{1},\widetilde{q}_{2},\widetilde{p}_{1}^{(01)},\widetilde{p}_{2}^{(01)}\right)$ is the same as system (4.1) except changing the parameters and has the accessible singularity
$\left\{\widetilde{q}_{1}=\widetilde{q}_{2}=\widetilde{p}_{1}^{(01)}=0\right\}.$ (5.9)
It is obvious that point (5.8) is contained in accessible singularity (5.9). On the other hand, we have given the coordinate resolving
$C_{3}^{(01)}=\left\{q_{1}=q_{2}=p_{1}^{(01)}=0\right\}$
in Proposition 4.1. Hence we can resolve the accessible singularity (5.8) as
$\left(\widetilde{x}_{3}^{[01]},\widetilde{y}_{3}^{[01]},\widetilde{z}_{3}^{[01]},\widetilde{w}_{3}^{[01]}\right)=\left(-(\widetilde{q}_{1}\widetilde{p}_{1}+\widetilde{q}_{2}\widetilde{p}_{2}-\alpha_{3}-\alpha_{5}+\eta)\widetilde{p}_{1},\widetilde{q}_{2}\widetilde{p}_{1},\dfrac{1}{\widetilde{p}_{1}},\dfrac{\widetilde{p}_{2}}{\widetilde{p}_{1}}\right)$ (5.10)
and it is a symplectic transformation. Note that the system in terms of $\left(\widetilde{x}_{3}^{[01]},\widetilde{y}_{3}^{[01]},\widetilde{z}_{3}^{[01]},\widetilde{w}_{3}^{[01]}\right)$ is described as a polynomial Hamiltonian system.
In the last, we see the solution to the system in terms of $\left(\widetilde{x}_{3}^{[01]},\widetilde{y}_{3}^{[01]},\widetilde{z}_{3}^{[01]},\widetilde{w}_{3}^{[01]}\right)$. Relations (5.7) and (5.10) imply
$\begin{split}\left(\widetilde{x}_{3}^{[01]},\widetilde{y}_{3}^{[01]},\widetilde{z}_{3}^{[01]},\widetilde{w}_{3}^{[01]}\right)&=\left(\frac{-q_{1}+(\alpha_{5}-\eta)p_{1}^{(01)}-q_{2}p_{2}^{(01)}}{\left(p_{1}^{(01)}\right)^{2}},\frac{\alpha_{3}p_{1}^{(01)}+q_{2}p_{2}^{(01)}}{p_{1}^{(01)}p_{2}^{(01)}},p_{1}^{(01)},p_{2}^{(01)}\right).\end{split}$ (5.11)
Substituting solution (5.2) with (5.3), (5.4), (5.5) and (5.6) to relation (5.11), we obtain
$\widetilde{x}_{3}^{[01]}$ $=(A_{1}a_{2}+A_{2}d_{2}+A_{3})+O(t-t_{0}),$
$\widetilde{y}_{3}^{[01]}$ $=\frac{A_{4}d_{2}+A_{5}}{\alpha_{3}}+O(t-t_{0}),$
$\widetilde{z}_{3}^{[01]}$ $=O(t-t_{0}),$
$\widetilde{w}_{3}^{[01]}$ $=O(t-t_{0}),$
where $A_{1},\ldots,A_{5}\in\mathbb{C}(t_{0})[\alpha_{0},\ldots,\alpha_{5},\eta]$. Hence the solution to system (4.1) with initial condition (5.1) is transformed to that in terms of $\left(\widetilde{x}_{3}^{[01]},\widetilde{y}_{3}^{[01]},\widetilde{z}_{3}^{[01]},\widetilde{w}_{3}^{[01]}\right)$ whose initial condition is given by
$\left(\widetilde{x}_{3}^{[01]},\widetilde{y}_{3}^{[01]},\widetilde{z}_{3}^{[01]},\widetilde{w}_{3}^{[01]}\right)=\left(A_{1}a_{2}+A_{2}d_{2}+A_{3},\frac{A_{4}d_{2}+A_{5}}{\alpha_{3}},0,0\right).$
We can see that the arbitrary complex numbers $a_{2}$ and $d_{2}$ in solution (5.2) appear in this initial condition.
6 Main result
We have given 9 and 12 coordinates resolving accessible singularities (3.3) and (3.4) in Proposition 4.1 and 5.1 respectively. Hence we can construct an initial value space of system (1.1).
Theorem 6.1.
An initial value space of system (1.1) is constructed by glueing 24 coordinates with relations
$\left(q_{1}^{(1)},q_{2}^{(1)},p_{1}^{(1)},p_{2}^{(1)}\right)$ $=\left(\frac{1}{q_{1}},\frac{q_{2}}{q_{1}},-q_{1}(q_{1}p_{1}+q_{2}p_{2}+\eta),q_{1}p_{2}\right),$
$\left(q_{1}^{(2)},q_{2}^{(2)},p_{1}^{(2)},p_{2}^{(2)}\right)$ $=\left(\frac{q_{1}}{q_{2}},\frac{1}{q_{2}},q_{2}p_{1},-q_{2}(q_{1}p_{1}+q_{2}p_{2}+\eta)\right),$
$\left(x_{2}^{[01]},y_{2}^{[01]},z_{2}^{[01]},w_{2}^{[01]}\right)$ $=\left(-\{(q_{1}-1)p_{1}-\alpha_{0}\}p_{1},q_{2},\dfrac{1}{p_{1}},p_{2}\right),$
$\left(\widetilde{x}_{6}^{[01]},\widetilde{y}_{6}^{[01]},\widetilde{z}_{6}^{[01]},\widetilde{w}_{6}^{[01]}\right)$ $=r_{4}\left(x_{2}^{[01]},y_{2}^{[01]},z_{2}^{[01]},w_{2}^{[01]}\right),$
$\left(\widetilde{x}_{7}^{[01]},\widetilde{y}_{7}^{[01]},\widetilde{z}_{7}^{[01]},\widetilde{w}_{7}^{[01]}\right)$ $=r_{5}r_{4}\left(x_{2}^{[01]},y_{2}^{[01]},z_{2}^{[01]},w_{2}^{[01]}\right),$
$\left(x_{3}^{[01]},y_{3}^{[01]},z_{3}^{[01]},w_{3}^{[01]}\right)$ $=\left(-(q_{1}p_{1}+q_{2}p_{2}-\alpha_{5}+\eta)p_{1},q_{2}p_{1},\dfrac{1}{p_{1}},\dfrac{p_{2}}{p_{1}}\right),$
$\left(\widetilde{x}_{3}^{[01]},\widetilde{y}_{3}^{[01]},\widetilde{z}_{3}^{[01]},\widetilde{w}_{3}^{[01]}\right)$ $=r_{3}\left(x_{3}^{[01]},y_{3}^{[01]},z_{3}^{[01]},w_{3}^{[01]}\right),$
$\left(\widetilde{x}_{4}^{[01]},\widetilde{y}_{4}^{[01]},\widetilde{z}_{4}^{[01]},\widetilde{w}_{4}^{[01]}\right)$ $=r_{4}r_{3}\left(x_{3}^{[01]},y_{3}^{[01]},z_{3}^{[01]},w_{3}^{[01]}\right),$
$\left(x_{3}^{[02]},y_{3}^{[02]},z_{3}^{[02]},w_{3}^{[02]}\right)$ $=\left(q_{1}p_{2},-(q_{1}p_{1}+q_{2}p_{2}-\alpha_{5}+\eta)p_{2},\dfrac{p_{1}}{p_{2}},\dfrac{1}{p_{2}}\right),$
$\left(x_{3}^{[11]},y_{3}^{[11]},z_{3}^{[11]},w_{3}^{[11]}\right)$ $=\left(-\left(q_{1}^{(1)}p_{1}^{(1)}+q_{2}^{(1)}p_{2}^{(1)}-\alpha_{1}+\eta\right)p_{1}^{(1)},q_{2}^{(1)}p_{1}^{(1)},\dfrac{1}{p_{1}^{(1)}},\dfrac{p_{2}^{(1)}}{p_{1}^{(1)}}\right),$
$\left(x_{2}^{[12]},y_{2}^{[12]},z_{2}^{[12]},w_{2}^{[12]}\right)$ $=\left(q_{1}^{(1)},-\left\{\left(q_{2}^{(1)}-1\right)p_{2}^{(1)}-\alpha_{2}\right\}p_{2}^{(1)},p_{1}^{(1)},\dfrac{1}{p_{2}^{(1)}}\right),$
$\left(\widetilde{x}_{6}^{[12]},\widetilde{y}_{6}^{[12]},\widetilde{z}_{6}^{[12]},\widetilde{w}_{6}^{[12]}\right)$ $=r_{0}\left(x_{2}^{[12]},y_{2}^{[12]},z_{2}^{[12]},w_{2}^{[12]}\right),$
$\left(\widetilde{x}_{7}^{[12]},\widetilde{y}_{7}^{[12]},\widetilde{z}_{7}^{[12]},\widetilde{w}_{7}^{[12]}\right)$ $=r_{1}r_{0}\left(x_{2}^{[12]},y_{2}^{[12]},z_{2}^{[12]},w_{2}^{[12]}\right),$
$\left(x_{3}^{[12]},y_{3}^{[12]},z_{3}^{[12]},w_{3}^{[12]}\right)$ $=\left(q_{1}^{(1)}p_{2}^{(1)},-\left(q_{1}^{(1)}p_{1}^{(1)}+q_{2}^{(1)}p_{2}^{(1)}-\alpha_{1}+\eta\right)p_{2}^{(1)},\dfrac{p_{1}^{(1)}}{p_{2}^{(1)}},\dfrac{1}{p_{2}^{(1)}}\right),$
$\left(\widetilde{x}_{3}^{[12]},\widetilde{y}_{3}^{[12]},\widetilde{z}_{3}^{[12]},\widetilde{w}_{3}^{[12]}\right)$ $=r_{5}\left(x_{3}^{[12]},y_{3}^{[12]},z_{3}^{[12]},w_{3}^{[12]}\right),$
$\left(\widetilde{x}_{4}^{[12]},\widetilde{y}_{4}^{[12]},\widetilde{z}_{4}^{[12]},\widetilde{w}_{4}^{[12]}\right)$ $=r_{0}r_{5}\left(x_{3}^{[12]},y_{3}^{[12]},z_{3}^{[12]},w_{3}^{[12]}\right),$
$\left(x_{3}^{[21]},y_{3}^{[21]},z_{3}^{[21]},w_{3}^{[21]}\right)$ $=\left(-\left(q_{1}^{(2)}p_{1}^{(2)}+q_{2}^{(2)}p_{2}^{(2)}-\alpha_{3}+\eta\right)p_{1}^{(2)},q_{2}^{(2)}p_{1}^{(2)},\dfrac{1}{p_{1}^{(2)}},\dfrac{p_{2}^{(2)}}{p_{1}^{(2)}}\right),$
$\left(x_{2}^{[22]},y_{2}^{[22]},z_{2}^{[22]},w_{2}^{[22]}\right)$ $=\left(q_{1}^{(2)},-\left\{\left(q_{2}^{(2)}-\frac{1}{t}\right)p_{2}^{(2)}-\alpha_{4}\right\}p_{2}^{(2)},p_{1}^{(2)},\frac{1}{p_{2}^{(2)}}\right),$
$\left(\widetilde{x}_{6}^{[22]},\widetilde{y}_{6}^{[22]},\widetilde{z}_{6}^{[22]},\widetilde{w}_{6}^{[22]}\right)$ $=r_{2}\left(x_{2}^{[22]},y_{2}^{[22]},z_{2}^{[22]},w_{2}^{[22]}\right),$
$\left(\widetilde{x}_{7}^{[22]},\widetilde{y}_{7}^{[22]},\widetilde{z}_{7}^{[22]},\widetilde{w}_{7}^{[22]}\right)$ $=r_{3}r_{2}\left(x_{2}^{[22]},y_{2}^{[22]},z_{2}^{[22]},w_{2}^{[22]}\right),$
$\left(x_{3}^{[22]},y_{3}^{[22]},z_{3}^{[22]},w_{3}^{[22]}\right)$ $=\left(q_{1}^{(2)}p_{2}^{(2)},-\left(q_{1}^{(2)}p_{1}^{(2)}+q_{2}^{(2)}p_{2}^{(2)}-\alpha_{3}+\eta\right)p_{2}^{(2)},\dfrac{p_{1}^{(2)}}{p_{2}^{(2)}},\dfrac{1}{p_{2}^{(2)}}\right),$
$\left(\widetilde{x}_{3}^{[22]},\widetilde{y}_{3}^{[22]},\widetilde{z}_{3}^{[22]},\widetilde{w}_{3}^{[22]}\right)$ $=r_{1}\left(x_{3}^{[22]},y_{3}^{[22]},z_{3}^{[22]},w_{3}^{[22]}\right),$
$\left(\widetilde{x}_{4}^{[22]},\widetilde{y}_{4}^{[22]},\widetilde{z}_{4}^{[22]},\widetilde{w}_{4}^{[22]}\right)$ $=r_{2}r_{1}\left(x_{3}^{[22]},y_{3}^{[22]},z_{3}^{[22]},w_{3}^{[22]}\right).$
Remark 6.2.
As was seen in Section 1, Takenawa constructed another IVS of system (1.1) and gave a geometric interpretation in terms of the Néron-Severi bilattice in [20]. We have not clarified a relationship between this previous work and Theorem 6.1 yet, which is a future problem.
We also consider the holomorphy property of system (1.1). Choose 6 from 9 coordinates given in Proposition 4.1 and set
$\begin{split}(x_{0},y_{0},z_{0},w_{0})&=\left(x_{2}^{[01]},y_{2}^{[01]},z_{2}^{[01]},w_{2}^{[01]}\right)\\ &=\left(-\{(q_{1}-1)p_{1}-\alpha_{0}\}p_{1},q_{2},\dfrac{1}{p_{1}},p_{2}\right),\\ (x_{1},y_{1},z_{1},w_{1})&=\left(x_{3}^{[11]},y_{3}^{[11]},z_{3}^{[11]},w_{3}^{[11]}\right)\\ &=\left(-\left(q_{1}^{(1)}p_{1}^{(1)}+q_{2}^{(1)}p_{2}^{(1)}-\alpha_{1}+\eta\right)p_{1}^{(1)},q_{2}^{(1)}p_{1}^{(1)},\dfrac{1}{p_{1}^{(1)}},\dfrac{p_{2}^{(1)}}{p_{1}^{(1)}}\right),\\ (x_{2},y_{2},z_{2},w_{2})&=\left(x_{2}^{[12]},y_{2}^{[12]},z_{2}^{[12]},w_{2}^{[12]}\right)\\ &=\left(q_{1}^{(1)},p_{1}^{(1)},-\left\{\left(q_{2}^{(1)}-1\right)p_{2}^{(1)}-\alpha_{2}\right\}p_{2}^{(1)},\dfrac{1}{p_{2}^{(1)}}\right),\\ (x_{3},y_{3},z_{3},w_{3})&=\left(x_{3}^{[21]},y_{3}^{[21]},z_{3}^{[21]},w_{3}^{[21]}\right)\\ &=\left(-\left(q_{1}^{(2)}p_{1}^{(2)}+q_{2}^{(2)}p_{2}^{(2)}-\alpha_{3}+\eta\right)p_{1}^{(2)},q_{2}^{(2)}p_{1}^{(2)},\dfrac{1}{p_{1}^{(2)}},\dfrac{p_{2}^{(2)}}{p_{1}^{(2)}}\right),\\ (x_{4},y_{4},z_{4},w_{4})&=\left(x_{2}^{[22]},y_{2}^{[22]},z_{2}^{[22]},w_{2}^{[22]}\right)\\ &=\left(q_{1}^{(2)},-\left\{\left(q_{2}^{(2)}-\frac{1}{t}\right)p_{2}^{(2)}-\alpha_{4}\right\},p_{1}^{(2)},\frac{1}{p_{2}^{(2)}}p_{2}^{(2)}\right),\\ (x_{5},y_{5},z_{5},w_{5})&=\left(x_{3}^{[01]},y_{3}^{[01]},z_{3}^{[01]},w_{3}^{[01]}\right)\\ &=\left(-(q_{1}p_{1}+q_{2}p_{2}-\alpha_{5}+\eta)p_{1},q_{2}p_{1},\dfrac{1}{p_{1}},\dfrac{p_{2}}{p_{1}}\right).\end{split}$ (6.1)
Then we obtain the following proposition.
Proposition 6.3.
There exists a unique polynomial Hamiltonian system of degree 5 with respect to $q_{1},q_{2},p_{1},p_{2}$ which is holomorphic on each coordinates $(x_{i},y_{i},z_{i},w_{i})\ (i=0,\ldots,5)$. Moreover, this Hamiltonian system coincides with (1.1).
Proof.
Let $K$ be a Hamiltonian of degree 5 described as
$K=\sum_{k=1}^{5}\sum_{\begin{subarray}{c}j_{1}+j_{2}+j_{3}+j_{4}=k,\\[1.0pt] j_{1}\geq 0,\,j_{2}\geq 0,\,j_{3}\geq 0,\,j_{4}\geq 0\end{subarray}}a_{j_{1},j_{2},j_{3},j_{4}}q_{1}^{j_{1}}q_{2}^{j_{2}}p_{1}^{j_{3}}p_{2}^{j_{4}}.$ (6.2)
We first consider the coordinate $(x_{0},y_{0},z_{0},w_{0})$. The first relation of (6.1) is rewritten to
$q_{1}=1+(\alpha_{0}-x_{0}z_{0})z_{0},\quad q_{2}=y_{0},\quad p_{1}=\frac{1}{z_{0}},\quad p_{2}=w_{0}.$ (6.3)
Substituting relation (6.3) to (6.2), we obtain
$K=\frac{a_{0,0,5,0}}{z_{0}^{5}}+\frac{F_{1}}{z_{0}^{4}},$
where $F_{1}\in\mathbb{C}[x_{0},y_{0},z_{0},w_{0}]$. Then the coefficient $a_{0,0,5,0}$ is determined uniquely as $a_{0,0,5,0}=0$ and the Hamiltonian $K$ is descried as
$K=\frac{a_{0,0,4,0}+a_{1,0,4,0}+a_{0,1,4,0}y_{0}+a_{0,0,4,1}w_{0}}{z_{0}^{4}}+\frac{F_{2}}{z_{0}^{3}},$
where $F_{2}\in\mathbb{C}[x_{0},y_{0},z_{0},w_{0}]$. It follows that
$a_{0,0,4,0}+a_{1,0,4,0}=0,\quad a_{0,1,4,0}=0,\quad a_{0,0,4,1}=0.$
We repeat this procedure until $K$ is a polynomial with respect to $z_{0}$ and do the same for the other coordinates $(x_{i},y_{i},z_{i},w_{i})\ (i=1,\ldots,5)$. Then all of the coefficients $a_{j_{1},j_{2},j_{3},j_{4}}$ are determined uniquely and the Hamiltonian $K$ coincides with that of system (1.1). Note that we have to consider the Hamiltonian $K+\frac{p_{2}^{(2)}}{t^{2}}$ for the coordinate $(x_{4},y_{4},z_{4},w_{4})$ because the independent variable $t$ appears in the fifth relation of (6.1). ∎
Remark 6.4.
Sasano considered the holomorphy property of system (1.1) in terms of 7 coordinates in [13]. The number of coordinates in Proposition 6.3 is 1 less than that in [13] because we always assume that $\alpha_{0}+\cdots+\alpha_{5}=1$.
Appendix A Proof of Lemma 3.1
We prove only for $W_{01}\times B$. System (2.6) implies that the accessible singularities on $W_{01}\times B$ are given by
$p_{1}^{(01)}=F_{1}^{(01)}=F_{2}^{(01)}=G_{2}^{(01)}=0.$
We set
$\widetilde{F}_{1}^{(01)}$ $=F_{1}^{(01)}\big{|}_{p_{1}^{(01)}=0}$
$=(q_{1}-1)\left\{2q_{1}(q_{1}-t)+(q_{1}+q_{2})(q_{2}-t)p_{2}^{(01)}\right\},$
$\widetilde{F}_{2}^{(01)}$ $=F_{2}^{(01)}\big{|}_{p_{1}^{(01)}=0}$
$=(q_{2}-t)\left\{(q_{1}-1)(q_{1}+q_{2})+2q_{2}(q_{2}-1)p_{2}^{(01)}\right\},$
$\widetilde{G}_{2}^{(01)}$ $=G_{2}^{(01)}\big{|}_{p_{1}^{(01)}=0}$
$=\left[\left\{2q_{1}^{2}-2q_{1}q_{2}-(t+1)q_{1}+2q_{2}\right\}+\left\{2q_{1}q_{2}-2q_{2}^{2}-2tq_{1}+(t+1)q_{2}\right\}p_{2}^{(01)}\right]p_{2}^{(01)}.$
Then the accessible singularities are given by a system of algebraic equations
$\widetilde{F}_{1}^{(01)}=\widetilde{F}_{2}^{(01)}=\widetilde{G}_{2}^{(01)}=0.$ (A.1)
In the following, we solve it for $q_{1}$, $q_{2}$ and $p_{2}^{(01)}$.
We first assume that $q_{1}-1=0$. Under this condition, system (A.1) is described as
$q_{2}(q_{2}-1)(q_{2}-t)p_{2}^{(01)}=\left[t-1+\left\{2q_{2}^{2}-(t+3)q_{2}+2t\right\}p_{2}^{(01)}\right]p_{2}^{(01)}=0.$
Hence we obtain the solutions of (A.1) as
$q_{1}-1=p_{2}^{(02)}=0,$ (A.2)
$q_{1}-1=q_{2}=2tp_{2}^{(02)}+t-1=0,$ (A.3)
$q_{1}-1=q_{2}-1=p_{2}^{(02)}+1=0,$ (A.4)
$q_{1}-1=q_{2}-t=tp_{2}^{(02)}+1=0.$ (A.5)
We next assume that $q_{1}-1\neq 0$ and $q_{2}-t=0$. Under this condition, the system $\frac{\widetilde{F}_{1}^{(01)}}{q_{1}-1}=\widetilde{F}_{2}^{(01)}=\widetilde{G}_{2}^{(01)}=0$ is described as
$q_{1}(q_{1}-t)=\left[\left\{2q_{1}^{2}-(3t+1)q_{1}+2t\right\}-t(t-1)p_{2}^{(01)}\right]p_{2}^{(01)}=0.$
Hence we obtain the solutions of (A.1) as
$q_{1}=q_{2}-t=p_{2}^{(01)}=0,$ (A.6)
$q_{1}=q_{2}-t=(t-1)p_{2}^{(01)}-2=0,$ (A.7)
$q_{1}-t=q_{2}-t=p_{2}^{(01)}=0,$ (A.8)
$q_{1}-t=q_{2}-t=p_{2}^{(01)}+1=0.$ (A.9)
We next assume that $q_{1}-1\neq 0$, $q_{2}-t\neq 0$ and $p_{2}^{(01)}=0$. Under this condition, the system $\frac{\widetilde{F}_{1}^{(01)}}{q_{1}-1}=\frac{\widetilde{F}_{2}^{(01)}}{q_{2}-t}=\widetilde{G}_{2}^{(01)}=0$ is described as
$q_{1}(q_{1}-t)=q_{1}+q_{2}=0.$
Hence we obtain the solutions of (A.1) as
$q_{1}=q_{2}=p_{2}^{(01)}=0,$ (A.10)
$q_{1}-t=q_{2}+t=p_{2}^{(01)}=0.$ (A.11)
In the last, we assume that $q_{1}-1\neq 0$, $q_{2}-t\neq 0$ and $p_{2}^{(01)}\neq 0$. Then the system $\frac{\widetilde{F}_{1}^{(01)}}{q_{1}-1}=\frac{\widetilde{F}_{2}^{(01)}}{q_{2}-t}=\frac{\widetilde{G}_{2}^{(01)}}{p_{2}^{(01)}}=0$ is described as
$A\begin{pmatrix}1\\ p_{2}^{(01)}\end{pmatrix}=\begin{pmatrix}0\\ 0\end{pmatrix},\quad A=\begin{pmatrix}2q_{1}(q_{1}-t)&(q_{1}+q_{2})(q_{2}-t)\\ (q_{1}-1)(q_{1}+q_{2})&2q_{2}(q_{2}-1)\\ 2q_{1}^{2}-2q_{1}q_{2}-(t+1)q_{1}+2q_{2}&2q_{1}q_{2}-2q_{2}^{2}-2tq_{1}+(t+1)q_{2}\end{pmatrix}$. (A.12)
It is necessary that all of the minor determinants of the matrix $A$ are zero, namely
$(q_{1}-q_{2})X=(q_{1}-q_{2})Y=(q_{1}-q_{2})Z=0,$
where
$X=q_{1}^{2}q_{2}-q_{1}q_{2}^{2}-tq_{1}^{2}-3(t-1)q_{1}q_{2}+q_{2}^{2}+tq_{1}-tq_{2},$
$Y=2q_{1}^{2}q_{2}-2q_{1}q_{2}^{2}-2tq_{1}^{2}-3(t-1)q_{1}q_{2}+2q_{2}^{2}+2tq_{1}+(t-3)q_{2},$
$Z=-2q_{1}^{2}q_{2}+2q_{1}q_{2}^{2}+2tq_{1}^{2}+3(t-1)q_{1}q_{2}-2q_{2}^{2}-t(3t-1)q_{1}+2tq_{2}.$
We first consider the case $q_{1}-q_{2}=0$. Then the matrix $A$ is described as
$A=\begin{pmatrix}2q_{1}(q_{1}-t)&2q_{1}(q_{1}-t)\\ 2q_{1}(q_{1}-1)&2q_{1}(q_{1}-1)\\ -(t-1)q_{1}&-(t-1)q_{1}\end{pmatrix}$,
from which we obtain two solutions $q_{1}=q_{2}=0$ and $p_{2}^{(01)}+1=0$ for (A.12). We next consider the case $q_{1}-q_{2}\neq 0$ and $X=Y=Z=0$. Since $Y+Z=-3(t-1)(tq_{1}-q_{2})=0$, we have
$X=t(t-1)q_{1}(q_{1}+1)^{2},\quad Y=t(t-1)q_{1}(q_{1}+1)(2q_{1}-1),\quad Z=-t(t-1)q_{1}(q_{1}+1)(2q_{1}-1).$
It implies
$q_{1}+1=q_{2}+t=0.$
Then the matrix $A$ is described as
$A=\begin{pmatrix}2(t+1)&2t(t+1)\\ 2(t+1)&2t(t+1)\\ -3(t-1)&-3t(t-1)\end{pmatrix}$,
from which we obtain a solution $tp_{2}^{(01)}+1=0$ for (A.12). Hence we obtain three solutions of (A.1) as
$q_{1}=q_{2}=0,\quad p_{2}^{(01)}\neq 0,$ (A.13)
$q_{1}-q_{2}=p_{2}^{(01)}+1=0,$ (A.14)
$q_{1}+1=q_{2}+t=tp_{2}^{(01)}+1=0.$ (A.15)
We have obtained 13 solutions. Among them, solutions (A.4) and (A.9) are contained in (A.14). Moreover, (A.10) and (A.13) can be summarized. Therefore the accessible singularities $C_{1}^{(01)},\ldots,C_{3}^{(01)}$ are derived from (A.14), (A.2), (A.10) and (A.13) and $P_{1}^{(01)},\ldots,P_{7}^{(01)}$ from (A.8), (A.11), (A.6), (A.7), (A.3), (A.5), (A.15).
We can prove for the other $W_{ij}\times B$ in a similar manner.
Acknowledgement
This work was supported by JSPS KAKENHI Grant Number 20K03645.